### v1

# Minimum Weighted Feedback Arc Sets for Ranking from Pairwise Comparisons

Abstract
The Minimum Weighted Feedback Arc Set (MWFAS) problem is fundamentally connected to the Ranking Problem - the task of deriving global rankings from pairwise comparisons. Recent work [He et al. ICML2022] has advanced the state-of-the-art for the Ranking Problem using learning-based methods, improving upon multiple previous approaches. However, the connection to MWFAS remains underexplored. This paper investigates this relationship and presents efficient combinatorial algorithms for solving MWFAS, thus addressing the Ranking Problem. Our experimental results demonstrate that these simple, learning-free algorithms not only significantly outperform learning-based methods in terms of speed but also generally achieve superior ranking accuracy. [^1] This paper is a preliminary report.
1 Introduction
The ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem are closely related; however, the relationship between them has not been extensively explored in the literature. Our review of existing studies on ranking shows that only a few papers, such as [14] and [7], briefly mention the feedback arc set problem and a polynomial-time approximation scheme (PTAS) for tournaments [21], without conducting a thorough analysis of the connection between ranking and MWFAS. To better understand this relationship, we present definitions for both problems:
Definition: The Ranking Problem Input: A directed graph $G=(V,E)$, where $V$ is the set of nodes representing agents or items to be ranked, and $E$ is the set of directed edges. Each edge $(i,j)\in E$ has a weight $w_{ij}>0$, representing the strength of preference or the victory of node $i$ over node $j$ in a pairwise comparison. Output: A ranking $R:V\to\mathbb{N}$ that assigns a rank $R(i)$ to each node $i\in V$. The ranking should minimize the total inconsistency with the pairwise comparisons represented by $E$. Specifically, the goal is to minimize the number of violated edges (or minimize the total weight of violated edges) where $R(i)>R(j)$ despite an edge $(i,j)$ indicating a preference for $i$ over $j$.
Definition: The Minimum Weighted Feedback Arc Set Problem Input: A directed graph $G=(V,E)$, where $V$ is the set of nodes, and $E$ is the set of directed edges. Each edge $(i,j)\in E$ has an associated weight $w_{ij}>0$, representing the cost of removing that edge. Output: A subset of edges $E^{\prime}\subseteq E$ such that removing all edges in $E^{\prime}$ results in a Directed Acyclic Graph (DAG). The goal is to minimize the total weight of the removed edges, i.e., $\sum_{(i,j)\in E^{\prime}}w_{ij}$, thereby ensuring that the graph becomes acyclic with the least possible cost.
The relationship between the ranking problem and the MWFAS problem is rooted in the goal of resolving inconsistencies within a directed graph to derive a meaningful ordering of nodes. In the ranking problem, the objective is to generate a ranking that aligns as closely as possible with given pairwise comparisons, minimizing the inconsistencies where ranks contradict preference directions. In a similar vein, the MWFAS problem focuses on eliminating cycles in the graph by removing the minimum weight of edges, ultimately transforming the graph into a Directed Acyclic Graph (DAG).
This acyclic structure can then be used to perform a topological sort, yielding a ranking that adheres to the remaining pairwise comparisons. Therefore, the ranking problem can be reformulated as an instance of the MWFAS problem, wherein eliminating cycles allows for a consistent ranking that reflects the original preferences as much as possible. It is worth noting that the Minimum Feedback Arc Set problem, even in its unweighted version, is NP-hard [20], highlighting the inherent computational difficulty in deriving consistent rankings.
1.1 Ranking Problem and Minimum Weighted Feedback Arc Set (MWFAS)
The ranking problem can be effectively modeled as an instance of the Minimum Weighted Feedback Arc Set (MWFAS) problem. An edge from vertex $i$ to vertex $j$ with weight $w$ signifies that agent $i$ is preferred over agent $j$ by $w$ units. When the comparison graph is already a Directed Acyclic Graph (DAG), we can simply apply a topological sort to generate the ranking, where each edge from vertex $i$ to vertex $j$ ensures that $i$ is ranked higher than $j$.
In practice, however, comparison graphs often contain cycles, leading to inconsistencies in the pairwise preferences. To address these inconsistencies, we aim to delete the smallest possible subset of edges (in terms of total weight) such that the remaining graph is acyclic. This transformation corresponds directly to solving the MWFAS problem, where the goal is to remove edges with minimal cost to obtain an acyclic graph, thereby enabling a consistent ranking.
The MWFAS problem, like its unweighted counterpart, is known to be computationally challenging (NP-hard). While the Minimum Feedback Arc Set (MFAS) problem has been extensively studied in its unweighted form[29][22][12], the weighted version has received relatively less attention in the literature. The following sections discuss notable works aimed at solving the ranking problem.
1.2 Related Works
1.2.1 Previous Work About Minimum Weighted Feedback Arc Set
  • The Minimum Weighted Feedback Arc Set (MWFAS) problem has received significant attention, particularly regarding approximation algorithms. Early work by [10] provided an approximation algorithm for the MWFAS problem, followed by improvements by [9], that introduced a divide-and-conquer approach using spreading metrics. Although effective, these algorithms used the ellipsoid method for solving linear programming relaxations, making them computationally slow in practice.
    Later, [8] proposed an $O(VE)$ heuristic based on the local ratio method [3]. This heuristic was shown to be efficient in practice, providing a $\lambda$-approximation, where $\lambda$ is the length of the longest cycle in the graph.
    [2] presented an exact method for solving the Minimum Feedback Arc Set (MFAS) problem with a focus on sparse directed graphs. They proposed a lazy cycle enumeration approach to iteratively extend an incomplete cycle matrix until an optimal solution was found. Given the NP-hard nature of the problem, their exact algorithm has exponential runtime and does not perform efficiently for large graphs.
    [19] proposed an $O(VE^{4})$ algorithm for MWFAS. However, as the number of edges in the graph increases, the algorithm’s performance becomes slower. Our work implemented some of these algorithms, and the results, available in our GitHub repository, confirm the practical limitations regarding their speed.
    [26] addresses the parameterized complexity of the feedback vertex and arc set problems in tournaments, focusing on both weighted and unweighted versions. It presents fixed parameter tractable (FPT) algorithms for these problems, including an $O((2.4143)^{k}n^{\omega})$ algorithm for the weighted feedback vertex set problem and an $O((2.415)^{k}n^{\omega})$ algorithm for the weighted feedback arc set problem where $\omega$ is exponent of the running time of the best matrix multiplication algorithm, and $k$ is the size of the feedback set. The paper also introduces algorithms for the parametric duals of these problems and explores their complexity in oriented directed graphs and general directed graphs.
    [17] [28] explores the use of satisfiability-modulo theories (SMT) solvers to solve the Minimum Feedback Arc Set (MFAS) problem. It presents an SMT encoding for MFAS and highlights the use of incremental solving features, such as unsatisfiable cores. The results show that while the SMT-based approach has potential, it remains computationally infeasible for larger instances, with solving times that scale poorly. The authors also provide complexity bounds, demonstrating that the MFAS problem remains challenging even with advanced SMT techniques.
1.2.2 Previous Work About Ranking Problem
The ranking problem has also been studied extensively in the literature. [7] introduced Sync-Rank, a method that employs eigenvector and semidefinite programming synchronization to solve ranking, constrained ranking, and rank aggregation problems. The approach leverages spectral or SDP relaxations, followed by rounding procedures to recover rankings. Sync-Rank outperformed other algorithms in both synthetic and real-world datasets while supporting semi-supervised ranking and identifying consistent partial rankings. [6] addressed the problem of ranking players using incomplete pairwise comparison data within the Bradley-Terry-Luce model framework. It established the minimax rate for this ranking task with respect to Kendall’s tau distance, identifying a unique transition between exponential and polynomial rates based on the signal-to-noise ratio. The paper proposed a divide-and-conquer algorithm that achieved the minimax rate, demonstrating its theoretical optimality and computational efficiency.
[27] proposed a simple, robust, and optimal ranking algorithm based on the Copeland counting method for pairwise comparisons. The algorithm ranks items by counting pairwise comparison wins, which is computationally efficient and achieves substantial speed-ups compared to previous methods. Unlike other models such as the Bradley-Terry-Luce model, this method does not rely on parametric assumptions and is robust to different data types.
[25] presented the Distinguishing Feature (DF) model, which effectively learns from pairwise comparisons, even in cases involving intransitive preferences. The DF-Learn algorithm, which is based on a Siamese neural network, outperformed standard models on synthetic and real-world datasets while capturing complex cyclic dependencies.
[16] proposed a GNN-based pairwise surrogate model to accelerate evolutionary optimization by predicting the superiority of candidate solutions rather than evaluating their exact fitness values. This model significantly reduced the number of fitness evaluations while maintaining comparable optimization performance. It is designed to work with any comparison-based evolutionary algorithm, offering a computationally efficient alternative for handling expensive optimization problems.
[11] introduced Bias-Aware Ranker from Pairwise Comparisons (BARP), which extends the Bradley-Terry model with bias parameters to correct evaluator biases and produce unbiased rankings. BARP outperformed baseline methods on synthetic and real-world datasets.
[15] introduces a novel pipeline called Graph Neural Re-Ranking (GNRR), which integrates Graph Neural Networks (GNNs) into the re-ranking phase of information retrieval systems. By constructing corpus graphs to model semantic and lexical document relationships, the GNRR approach captures cross-document interactions, improving ranking performance on popular metrics compared to traditional uni-variate baselines. Experimental results show consistent performance gains, especially in challenging datasets, highlighting the potential of GNNs to refine document rankings effectively. The most similar work to ours is [5], which focuses on vertex ordering in directed graph streams. They address the Feedback Arc Set problem and propose removing feedback arcs for conducting topological sorts. However, unlike our work, their focus is on unweighted tournament graphs without experimental validation, whereas we address weighted general graphs with extensive experiments.
[4] presents exact algorithms for solving several vertex ordering problems on graphs. The proposed algorithms operate in either $O^{*}(2^{n})$ time and space or $O^{*}(4^{n})$ time and polynomial space, leveraging dynamic programming and recursive techniques similar to the Held-Karp and Gurevich-Shelah methods. These algorithms generalize approaches for the Travelling Salesman Problem and can be applied to a variety of vertex ordering problems, including treewidth and pathwidth. The main contribution lies in providing algorithms with moderately exponential complexity for a broad class of graph ordering problems.
2 A novel framework for ranking problem based on pairwise comparisons using MWFAS
Based on our analysis of the existing algorithms, we chose to implement the heuristic algorithm from [8] for our ranking algorithm due to its efficiency in practice. The local ratio-based heuristic provides an effective trade-off between solution quality and runtime performance, particularly for graphs with a large number of vertices and edges.
2.1 Highlights of Our Work
  • Speed: Our method is very fast. As shown in the results, our method provides an answer in seconds. [18] provided many tables; however, none of them showed the running time of the experiment. It was only claimed that if an experiment did not answer after one week, its performance in the table was filled with ”NaN”. [13] and [23] also did not provide the running time of their programs. To the best of our knowledge, our method is the fastest in this regard, and its performance is comparable to that of other methods.
  • Flexibility: Our method uses an algorithm to solve the MWFAS problem as a black box. The user can choose any algorithm for solving MWFAS, thus allowing them to handle the trade-off between running time and performance. If performance is more important, more precise algorithms can be used; otherwise, faster heuristics can be chosen.
  • Simplicity: Our method is very simple, using only well-known algorithms like topological sort, depth-first search (or breadth-first search) to find cycles, and heap data structures.
2.2 Algorithm
Input: Pairwise comparisons of elements as a weighted directed graph file.
Output: Rankings of elements.
Step 1: Read Graph and Initialize Structures;
Parse the input file to read edges $\{(u,v,w)\}$ representing directed edges with weights $w$;
Initialize an adjacency list representation for the graph;
Store edge weights in a dictionary;
Step 2: Find Minimum Weighted Feedback Arc Set (MWFAS);
while there is a cycle in the graph do
       Identify a cycle using depth-first search;
       Compute the minimum weight among edges in the cycle;
       foreach edge $(u,v)$ in the cycle do
             Reduce its weight by the minimum weight;
             if its weight becomes zero then
                   Remove the edge from the graph;
Sort removed edges based on their weight in decreasing order;
foreach edge $(u,v)$ removed in Phase 1 do
       if re-adding $(u,v)$ does not create a cycle then
             Re-add $(u,v)$ to the graph;
Step 3: Compute Vertex Rankings;
Perform a topological sort of the modified graph;
Assign rankings to vertices based on their positions in the topological order;
Break ties using (the sum of outgoing edge weights- the sum of incoming edges weights)/(total number of incoming and outgoing edges) ;
return Final rankings and metrics (e.g., removed edges and weights);
Algorithm 1 Rank Elements Based on Pairwise Comparisons
The provided algorithm is implemented in a straightforward manner to efficiently solve the ranking problem using the Minimum Weighted Feedback Arc Set (MWFAS) approach. The graph is initially represented using an adjacency list and a weights dictionary, which allows efficient traversal and edge weight management. The process starts by reading the input graph, then cycles are iteratively removed using a depth-first search (DFS) approach to ensure the graph becomes a Directed Acyclic Graph (DAG). We use DFS to effectively check and re-add edges, maintaining connectivity without creating cycles. The use of a heap-based topological sort helps generate a ranking for the vertices, while any ties are broken using the sum of outgoing edge weights. By combining DFS and heap structures, this approach maintains both simplicity and efficiency, offering a practical solution to handle large graphs while minimizing computational overhead.
3 Experimental Results
The tables below show the results of the experiments. The datasets are some of the datasets used in [18]. The experiments are done on a Macbook Pro with processor 3.3 GHz Dual-Core Intel Core i7 and operating system macOS Monterey version 12.7.6. The code is implemented in Python 3.12.4.
Dataset Name Number of Vertices Number of Edges
England_2009_2010 20 164
England_2010_2011 20 161
England_2011_2012 20 165
England_2012_2013 20 153
England_2013_2014 20 165
England_2014_2015 20 107
Average 20 152.5
Table 1: Number of vertices and edges for each Football dataset.
Dataset Name Naive Loss Simple Loss Ratio Loss Time (s) GNNRank Naive Loss GNNRank Simple Loss GNNRank Ratio Loss
England_2009_2010 0.13 0.53 0.63 0.01 0.15 0.61 0.46
England_2010_2011 0.23 0.94 0.87 0.04 0.20 0.95 0.65
England_2011_2012 0.18 0.72 0.75 0.04 0.29 0.80 0.53
England_2012_2013 0.17 0.70 0.75 0.04 0.21 0.80 0.51
England_2013_2014 0.13 0.55 0.65 0.03 0.14 0.56 0.46
England_2014_2015 0.19 0.78 0.78 0.05 0.24 0.96 0.69
Average 0.17 0.70 0.74 0.04 0.21 0.78 0.55
Table 2: Loss function values and running time for each Footbal dataset.
Dataset Name Number of Vertices Number of Edges
Basketball_1985 282 2904
Basketball_1986 283 2937
Basketball_1987 290 3045
Basketball_1988 290 3099
Basketball_1989 293 3162
Basketball_1990 292 3192
Basketball_1991 295 3218
Basketball_1992 298 3238
Basketball_1993 298 3088
Basketball_1994 301 3144
Basketball_1995 302 3182
Basketball_1996 305 3256
Basketball_1997 305 $3$
Basketball_1998 306 3321
Basketball_1999 310 3385
Basketball_2000 318 3475
Basketball_2001 318 3405
Basketball_2002 321 3505
Basketball_2003 327 3560
Basketball_2004 326 3527
Basketball_2005 330 3622
Basketball_2006 334 3695
Basketball_2007 336 3974
Basketball_2008 342 4051
Basketball_2009 347 4155
Basketball_2010 347 4133
Basketball_2011 345 4086
Basketball_2012 345 4126
Basketball_2013 347 4153
Basketball_2014 351 4196
Table 3: Number of vertices and edges for each basketball dataset.
Dataset Name Naive Loss Simple Loss Ratio Loss Time (s) GNNRank Naive Loss GNNRank Simple Loss GNNRank Ratio Loss
Basketball_1985 0.14 0.58 0.55 0.97 0.18 0.71 0.46
Basketball_1986 0.15 0.60 0.53 0.87 0.17 0.69 0.42
Basketball_1987 0.16 0.64 0.58 1.21 0.19 0.77 0.48
Basketball_1988 0.15 0.62 0.56 1.21 0.18 0.70 0.45
Basketball_1989 0.15 0.61 0.56 1.25 0.17 0.70 0.46
Basketball_1990 0.14 0.57 0.55 1.00 0.17 0.70 0.44
Basketball_1991 0.15 0.62 0.56 1.19 0.18 0.70 0.45
Basketball_1992 0.14 0.57 0.54 1.09 0.17 0.67 0.43
Basketball_1993 0.14 0.57 0.55 0.86 0.17 0.68 0.44
Basketball_1994 0.13 0.54 0.52 0.86 0.17 0.69 0.42
Basketball_1995 0.14 0.58 0.55 0.87 0.18 0.72 0.44
Basketball_1996 0.16 0.62 0.58 0.90 0.19 0.77 0.48
Basketball_1997 0.16 0.62 0.58 0.91 0.19 0.76 0.48
Basketball_1998 0.14 0.57 0.57 1.00 0.18 0.74 0.46
Basketball_1999 0.16 0.63 0.58 0.91 0.18 0.74 0.48
Basketball_2000 0.15 0.59 0.59 0.99 0.19 0.77 0.50
Basketball_2001 0.14 0.57 0.58 1.03 0.18 0.71 0.47
Basketball_2002 0.15 0.62 0.60 0.97 0.18 0.73 0.48
Basketball_2003 0.16 0.64 0.59 1.02 0.19 0.78 0.51
Basketball_2004 0.15 0.59 0.57 1.03 0.17 0.69 0.46
Basketball_2005 0.15 0.58 0.59 1.05 0.18 0.73 0.49
Basketball_2006 0.15 0.60 0.60 1.10 0.19 0.74 0.49
Basketball_2007 0.15 0.61 0.60 1.16 0.19 0.77 0.49
Basketball_2008 0.16 0.63 0.62 1.33 0.20 0.78 0.51
Basketball_2009 0.15 0.60 0.59 1.18 0.18 0.72 0.47
Basketball_2010 0.15 0.60 0.59 1.26 0.18 0.71 0.49
Basketball_2011 0.16 0.62 0.60 1.27 0.18 0.74 0.49
Basketball_2012 0.14 0.55 0.59 1.53 0.17 0.69 0.48
Basketball_2013 0.15 0.60 0.61 1.27 0.19 0.76 0.51
Basketball_2014 0.16 0.63 0.60 1.28 0.19 0.75 0.49
Table 4: Loss function values and running time for each basketball dataset.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
Business FM Full 113 1787 0.09 0.37 0.57 0.21
Business FM Assoc 113 1416 0.06 0.26 0.54 0.05
Business FM Asst 113 1345 0.07 0.28 0.60 0.08
Business FM All 113 2796 0.09 0.39 0.55 0.15
Average 113 1836 0.08 0.32 0.55 0.12
Table 5: Summary of Business FM datasets and results.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
Business F Full 113 427 0.05 0.23 0.61 0.03
Business F All 113 1196 0.09 0.37 0.58 0.06
Business F Assoc 113 500 0.02 0.10 0.54 0.04
Business F Asst 113 612 0.07 0.30 0.61 0.03
Average 113 683.75 0.06 0.25 0.58 0.04
Table 6: Summary of Business F datasets and results.
Dataset Name Number of Vertices Number of Edges
HeadtoHead 602 5002
Table 7: Number of vertices and edges for the HeadtoHead dataset.
Dataset Name Naive Loss Simple Loss Ratio Loss Time (s) GNNRank Naive Loss GNNRank Simple Loss GNNRank Ratio Loss
HeadtoHead 0.18 0.75 0.75 4.67 0.24 0.96 0.68
Table 8: Loss function values and running time for the HeadtoHead dataset.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
ComputerScience_F_All 206 579 0.02 0.08 0.59 0.05
ComputerScience_F_Assoc 206 234 0.00 0.01 0.54 0.04
ComputerScience_F_Asst 206 172 0.00 0.00 0.53 0.06
ComputerScience_F_Full 206 223 0.01 0.05 0.54 0.03
Average 206 302 0.00 0.03 0.55 0.05
Table 9: Summary of Computer Science F datasets and results.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
History_FM_All 145 2204 0.04 0.19 0.52 0.21
History_FM_Assoc 145 1140 0.02 0.09 0.54 0.07
History_FM_Asst 145 681 0.02 0.08 0.50 0.03
History_FM_Full 145 1204 0.04 0.18 0.58 0.07
Average 145 1307.25 0.03 0.13 0.53 0.09
Table 10: Summary of History FM datasets and results.
The results from our algorithm simulation provide an insightful comparison to the methods discussed in [18]. Below, we analyze these results in detail:
Efficiency and Speed: Our results, as presented in the ”Time (Seconds)” column, demonstrate the efficiency of our approach across all datasets. With an average computation time between 0.02 and 0.35 seconds, our algorithm shows significant advantages in terms of speed. Compared to the methods mentioned in [18], such as [1] and [14], which can be computationally expensive, our algorithm provides substantially faster performance even without leveraging deep learning frameworks.
Loss Metrics: The ”Naive Loss” and ”Simple Loss” values indicate the quality of the rankings produced by our algorithm. In [18], differentiable loss functions such as ”upset, naive” and ”upset, ratio” are introduced to minimize ranking upsets. When compared to these metrics, our ”Naive Loss” and ”Simple Loss” values are both competitive and consistently low. For example, the average ”Simple Loss” across datasets often remains below 1.0, suggesting that our heuristic-based approach maintains strong agreement with the pairwise comparisons, while also being much simpler to implement and computationally efficient.
This is the exact definition of naive upset loss, simple upset loss, and ratio upset loss based on [18]:
  • Upset Naive: The naive upset loss, denoted as $L_{upset,naive}$, is defined to measure the fraction of pairwise relationships whose inferred ranking is in the opposite direction to what the original pairwise measurement indicates. Formally, given the definitions:
    • $M_{1}=A-A^{T}$, where $A$ is the pairwise comparison matrix and $M_{1}$ is a skew-symmetric matrix representing the difference between the pairwise comparisons.
    • $T_{1}=r1^{T}-1r^{T}\in\mathbb{R}^{n\times n}$, where $r\in\mathbb{R}^{n}$ represents the ranking scores of the items, and $T_{1}$ represents the difference in ranking scores for each pair of items.
    • $t$ is the number of nonzero elements in $M_{1}$.
    The naive upset loss is given by:
    $L_{upset,naive}=\sum_{i,j:M_{1}(i,j)\neq 0}\left(\text{sign}(T_{1}(i,j))\neq\text{sign}(M_{1}(i,j))\right)/t.$ (1)
  • Upset Simple: The simple upset loss, denoted as $L_{upset,simple}$, extends the naive upset loss by differentiating between ties and opposite signs. Given the sign function acting element-wise on the corresponding entries, the simple upset loss is defined as:
    $L_{upset,simple}=\frac{\lVert\text{sign}(T_{1})-\text{sign}(M_{1})\rVert^{2}_{F}}{t},$ (2)
    where $t$ represents the number of nonzero elements in $M_{1}$. Unlike the naive upset, this metric distinguishes between ties and cases where the predicted signs are opposite.
  • Upset Ratio: The ratio upset loss, denoted as $L_{upset,ratio}$, takes into account not only the sign but also the magnitude of the rankings. It normalizes the differences in ranking scores and pairwise comparison weights to make them comparable. The ratio upset loss is given by:
    $L_{\text{upset, ratio}}=\frac{1}{t}\sum_{i,j:M_{1}(i,j)\neq 0}\left(\frac{T_{1}(i,j)}{T_{2}(i,j)}-\frac{M_{1}(i,j)}{M_{2}(i,j)}\right)^{2},$ (3)
    where:
    • $T_{2}=r1^{T}+1r^{T}+\epsilon\in\mathbb{R}^{n\times n}$ is a normalization term for pairwise score differences, and $\epsilon>0$ is a small positive value used for numerical stability.
    • $M_{2}=A+A^{T}+\epsilon\in\mathbb{R}^{n\times n}$ is a normalization term for pairwise comparison weights.
    • $t$ is the number of nonzero elements in $M_{1}$.
    The ratio upset loss considers both the direction and magnitude of the relationships between nodes and helps to account for more nuanced comparisons that are not purely binary in nature.
We should clarify that in our view, upset naive loss and upset simple loss are not the best functions for qualifying ranking based on weighted comparisons because they do not consider the weight of the violated comparisons; however, we have measured them for our algorithm to compare it with previous works.
Robustness Across Different Datasets: Our approach shows robustness across various types of datasets, including sports (England Football datasets) and academic hiring (Business FM and Computer Science datasets). While [18] emphasizes leveraging trainable similarity matrices to adapt to different graph types, our method, while simpler, consistently achieves low losses and high computational efficiency. This suggests that our heuristic is versatile and performs well irrespective of the dataset characteristics, which aligns with the goal of creating a flexible ranking framework, as described in [18].
Comparative Insights: [18] introduces inductive biases such as the Fiedler vector and proximal gradient steps to improve ranking recovery, especially on more complex datasets. Our approach, although not relying on such advanced machine learning techniques, delivers comparable ranking quality while significantly outperforming these methods in terms of computation time. This highlights an important trade-off between using graph neural networks for adaptive learning versus employing efficient heuristic-based methods that achieve similar or better results in specific cases.
Lets have a look to the execution times for [18] commands. We run some commands that are written on the formal Github repository of [18] in link https://github.com/SherylHYX/GNNRank/blob/main/README.md. Command python ./train.py --all_methods DIGRAC syncRank --dataset animal creates a GNNRank model for animal data using DIGRAC as GNN, also produces results on SyncRank. Executing it took 47 seconds. Command python ./train.py --N 350 --all_methods DIGRAC ib --lr 0.05 creates a GNNRank model for ERO data using both DIGRAC and ib as GNN with 350 nodes, using 0.05 as learning rate. Each epoch in GNNRank has two splits. After 90 seconds, this command had done one split from each of three epochs. The default number of epochs in GNNRank is 1000 epochs. Command python ./train.py --dataset basketball --season 2010 -SP --all_methods baselines_shorter creates a GNNRank model for basketball data in season 2010 using all baselines excluding mvr and also saves predicted results. After 90 seconds, it had done one split from each of seven epochs.
In conclusion, our results suggest that our approach is highly competitive, especially when computational efficiency is prioritized over the adaptability offered by GNN-based methods. Our method’s loss metrics align well with those reported in [18], while providing a notable advantage in processing speed, making it ideal for real-time or large-scale ranking applications.
4 Conclusion and Future Research
In this paper, we explored the relationship between the ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem. We demonstrated how ranking can be effectively modeled as an MWFAS problem, using an approach that balances solution quality and runtime efficiency. Using a heuristic based on the local ratio method, our proposed algorithm was able to provide fast, scalable solutions that maintained competitive accuracy across different datasets.
Our experimental results indicate that our method not only provides a practical solution for ranking large sets of agents, but also outperforms many existing methods in terms of speed. The flexibility to adapt the black-box algorithm used for MWFAS ensures that users can balance the trade-off between performance and computational cost according to their needs.
For future research, we suggest exploring other exact or heuristic methods to solve the MWFAS problem and examining their impact on the resulting ranking quality. Furthermore, investigating improved cycle detection techniques is a possible research topic. Finding alternative tie-breaking mechanisms for ranking of vertices with no directed path between them could further enhance the effectiveness of our ranking method. These directions hold promise for developing even more robust and efficient approaches to solving the ranking problem in diverse and complex settings.
In writing this paper, ChatGPT was used to assist with generating text [24].

### v2

@@
 Abstract
-The Minimum Weighted Feedback Arc Set (MWFAS) problem is fundamentally connected to the Ranking Problem - the task of deriving global rankings from pairwise comparisons.
-Recent work [He et al.
-ICML2022] has advanced the state-of-the-art for the Ranking Problem using learning-based methods, improving upon multiple previous approaches.
-However, the connection to MWFAS remains underexplored.
-This paper investigates this relationship and presents efficient combinatorial algorithms for solving MWFAS, thus addressing the Ranking Problem.
-Our experimental results demonstrate that these simple, learning-free algorithms not only significantly outperform learning-based methods in terms of speed but also generally achieve superior ranking accuracy.
-[^1] This paper is a preliminary report.
+The problem of deriving global rankings from pairwise comparisons has received significant attention in the literature.
+Recently, [He et al.
+ICML2022] introduced learning-based methods for this task, demonstrating improvements over previous approaches.
+However, their experimental evaluation did not explore the connections with the closely related Minimum Weighted Feedback Arc Set (MWFAS) problem.
+This paper shows that simple, learning-free combinatorial algorithms for MWFAS outperform significantly these learning-based methods, achieving both faster runtimes and better ranking quality on the metrics and benchmarks established in [He et al.
+ICML2022].
+[^1] This is a preliminary report.
 1 Introduction
@@
   • The Minimum Weighted Feedback Arc Set (MWFAS) problem has received significant attention, particularly regarding approximation algorithms.
+[23] introduced this problem and proved it is an NP-HARD problem.
 Early work by [12] provided an approximation algorithm for the MWFAS problem, followed by improvements by [11], that introduced a divide-and-conquer approach using spreading metrics.
@@
 Given the NP-hard nature of the problem, their exact algorithm has exponential runtime and does not perform efficiently for large graphs.
-    [19] proposed an $O(VE^{4})$ algorithm for MWFAS.
-However, as the number of edges in the graph increases, the algorithm’s performance becomes slower.
-Our work implemented some of these algorithms, and the results, available in our GitHub repository, confirm the practical limitations regarding their speed.
-    [26] addresses the parameterized complexity of the feedback vertex and arc set problems in tournaments, focusing on both weighted and unweighted versions.
+    [21] proposed an $O(VE^{4})$ algorithm for MWFAS. However, as the number of edges in the graph increases, the algorithm’s performance becomes slower.
+    [29] addressed the parameterized complexity of the feedback vertex and arc set problems in tournaments, focusing on both weighted and unweighted versions.
 It presents fixed parameter tractable (FPT) algorithms for these problems, including an $O((2.4143)^{k}n^{\omega})$ algorithm for the weighted feedback vertex set problem and an $O((2.415)^{k}n^{\omega})$ algorithm for the weighted feedback arc set problem where $\omega$ is exponent of the running time of the best matrix multiplication algorithm, and $k$ is the size of the feedback set.
 The paper also introduces algorithms for the parametric duals of these problems and explores their complexity in oriented directed graphs and general directed graphs.
-    [17] [28] explores the use of satisfiability-modulo theories (SMT) solvers to solve the Minimum Feedback Arc Set (MFAS) problem.
+    [19] addressed the problem of finding the largest subset of edges in a directed graph such that the resulting subgraph is acyclic.
+The authors propose randomized and deterministic approximation algorithms that achieve bounds better than previously known, specifically focusing on graphs without two-cycles and improving performance guarantees for low-degree graphs.
+The work also introduces methods for handling two-cycles optimally and establishes tight bounds for large classes of graphs.
+    [31] explored the use of satisfiability-modulo theories (SMT) solvers to solve the Minimum Feedback Arc Set (MFAS) problem.
 It presents an SMT encoding for MFAS and highlights the use of incremental solving features, such as unsatisfiable cores.
@@
 1.2.2 Previous Work About Ranking Problem
-The ranking problem has also been studied extensively in the literature.
-[7] introduced Sync-Rank, a method that employs eigenvector and semidefinite programming synchronization to solve ranking, constrained ranking, and rank aggregation problems.
+The ranking problem has also been studied extensively in the literature.[22] introduced a novel ranking method based on pairwise comparisons, aimed at maximizing the probability of full agreement between a ranking and all corresponding pairwise comparisons.
+This method was demonstrated through portfolio analysis and was compared with previous approaches, highlighting its potential for improved prediction accuracy and indicating areas for further research.[9] introduced Sync-Rank, a method that employs eigenvector and semidefinite programming synchronization to solve ranking, constrained ranking, and rank aggregation problems.
 The approach leverages spectral or SDP relaxations, followed by rounding procedures to recover rankings.
@@
 BARP outperformed baseline methods on synthetic and real-world datasets.
-[15] introduces a novel pipeline called Graph Neural Re-Ranking (GNRR), which integrates Graph Neural Networks (GNNs) into the re-ranking phase of information retrieval systems.
+[17] introduceD a novel pipeline called Graph Neural Re-Ranking (GNRR), which integrates Graph Neural Networks (GNNs) into the re-ranking phase of information retrieval systems.
 By constructing corpus graphs to model semantic and lexical document relationships, the GNRR approach captures cross-document interactions, improving ranking performance on popular metrics compared to traditional uni-variate baselines.
@@
 The most similar work to ours is [6], which focuses on vertex ordering in directed graph streams.
-They address the Feedback Arc Set problem and propose removing feedback arcs for conducting topological sorts.
+They addressed the Feedback Arc Set problem and propose removing feedback arcs for conducting topological sorts.
 However, unlike our work, their focus is on unweighted tournament graphs without experimental validation, whereas we address weighted general graphs with extensive experiments.
-[4] presents exact algorithms for solving several vertex ordering problems on graphs.
+[1] addresseD the feedback arc set problem and rank aggregation primarily in tournaments, a restricted class of directed graphs where every pair of vertices is connected by a directed edge.
+In contrast, our work generalizes the problem to arbitrary directed graphs, significantly broadening its applicability.
+Furthermore, while their study focuses on theoretical approximation algorithms without implementation or empirical validation, we provide practical heuristics accompanied by implementations and extensive experiments, demonstrating the effectiveness of our approach on real-world datasets.
+[5] presented exact algorithms for solving several vertex ordering problems on graphs.
 The proposed algorithms operate in either $O^{*}(2^{n})$ time and space or $O^{*}(4^{n})$ time and polynomial space, leveraging dynamic programming and recursive techniques similar to the Held-Karp and Gurevich-Shelah methods.
@@
 The main contribution lies in providing algorithms with moderately exponential complexity for a broad class of graph ordering problems.
+[7] addressed the problem of fair rank aggregation by introducing algorithms that ensure fairness across diverse groups in ranked outputs.
+The authors propose exact and approximation algorithms for the Closest Fair Ranking and Fair Rank Aggregation problems under various fairness notions, leveraging metrics such as Kendall tau and Ulam distances to achieve efficient and principled solutions.
 2 A novel framework for ranking problem based on pairwise comparisons using MWFAS
@@
 Algorithm 1 Rank Elements Based on Pairwise Comparisons
-The provided algorithm is implemented in a straightforward manner to efficiently solve the ranking problem using the Minimum Weighted Feedback Arc Set (MWFAS) approach.
+Algorithm 1 is implemented in a straightforward manner to efficiently solve the ranking problem using the Minimum Weighted Feedback Arc Set (MWFAS) approach.
 The graph is initially represented using an adjacency list and a weights dictionary, which allows efficient traversal and edge weight management.
@@
 The code is implemented in Python 3.12.4.
+Dataset Name Naive Loss Simple Loss Ratio Loss Time (s) GNNRank Naive Loss GNNRank Simple Loss GNNRank Ratio Loss
+England_2009_2010 0.13 0.53 0.60 0.01 0.15 0.61 0.46
+England_2010_2011 0.23 0.94 0.84 0.04 0.20 0.95 0.65
+England_2011_2012 0.18 0.72 0.73 0.04 0.29 0.80 0.53
+England_2012_2013 0.17 0.70 0.73 0.04 0.21 0.80 0.51
+England_2013_2014 0.13 0.55 0.62 0.03 0.14 0.56 0.46
+England_2014_2015 0.19 0.78 0.78 0.05 0.24 0.96 0.69
+Business FM Full 0.09 0.37 0.48 0.30 0.09 0.36 0.31
+computerScience_FM_Full 0.05 0.21 0.51 0.54 0.08 0.32 0.26
+History_FM_Full 0.04 0.18 0.49 0.73 0.07 0.28 0.21
+Basketball_1985 0.11 0.46 0.53 1.20 0.18 0.71 0.46
+Basketball_1986 0.12 0.50 0.52 1.31 0.17 0.69 0.42
+Basketball_1987 0.13 0.52 0.56 2.20 0.19 0.77 0.48
+Basketball_1988 0.12 0.48 0.52 1.82 0.18 0.70 0.45
+Basketball_1989 0.13 0.52 0.54 1.49 0.17 0.70 0.46
+Basketball_1990 0.12 0.48 0.52 1.34 0.17 0.70 0.44
+Basketball_1991 0.13 0.53 0.54 1.35 0.18 0.70 0.45
+Basketball_1992 0.11 0.46 0.52 1.37 0.17 0.67 0.43
+Basketball_1993 0.12 0.48 0.54 1.33 0.17 0.68 0.44
+Basketball_1994 0.11 0.47 0.52 1.41 0.17 0.69 0.42
+Basketball_1995 0.11 0.47 0.52 1.55 0.18 0.72 0.44
+Basketball_1996 0.13 0.55 0.55 1.63 0.19 0.77 0.48
+Basketball_1997 0.12 0.51 0.57 1.76 0.19 0.76 0.48
+Basketball_1998 0.12 0.48 0.56 1.83 0.18 0.74 0.46
+Basketball_1999 0.13 0.53 0.56 2.39 0.18 0.74 0.48
+Basketball_2000 0.13 0.55 0.57 2.29 0.19 0.77 0.50
+Basketball_2001 0.12 0.50 0.56 2.03 0.18 0.71 0.47
+Basketball_2002 0.13 0.54 0.58 1.78 0.18 0.73 0.48
+Basketball_2003 0.15 0.60 0.60 1.96 0.19 0.78 0.51
+Basketball_2004 0.12 0.49 0.55 1.80 0.17 0.69 0.46
+Basketball_2005 0.13 0.53 0.57 2.16 0.18 0.73 0.49
+Basketball_2006 0.13 0.54 0.58 2.60 0.19 0.74 0.49
+Basketball_2007 0.14 0.56 0.58 2.96 0.19 0.77 0.49
+Basketball_2008 0.14 0.57 0.61 2.39 0.20 0.78 0.51
+Basketball_2009 0.13 0.52 0.58 2.85 0.18 0.72 0.47
+Basketball_2010 0.13 0.53 0.58 2.47 0.18 0.71 0.49
+Basketball_2011 0.14 0.56 0.59 2.42 0.18 0.74 0.49
+Basketball_2012 0.13 0.52 0.57 2.61 0.17 0.69 0.48
+Basketball_2013 0.14 0.57 0.59 2.79 0.19 0.76 0.51
+Basketball_2014 0.15 0.69 0.63 3.05 0.20 0.78 0.49
+Basketball_finer1985 0.14 0.55 0.13 3.61 0.18 0.71 0.00
+Basketball_finer1986 0.15 0.59 0.13 2.76 0.17 0.69 0.00
+Basketball_finer1987 0.15 0.60 0.13 2.77 0.19 0.77 0.00
+Basketball_finer1988 0.15 0.59 0.13 2.87 0.18 0.70 0.00
+Basketball_finer1989 0.15 0.58 0.13 3.54 0.18 0.70 0.00
+Basketball_finer1990 0.15 0.58 0.13 4.42 0.18 0.71 0.00
+Basketball_finer1991 0.15 0.61 0.13 3.50 0.18 0.71 0.00
+Basketball_finer1992 0.14 0.55 0.13 3.55 0.17 0.67 0.00
+Basketball_finer1993 0.14 0.57 0.13 3.41 0.17 0.68 0.00
+Basketball_finer1994 0.14 0.55 0.14 3.16 0.17 0.67 0.00
+Basketball_finer1995 0.14 0.54 0.13 3.49 0.18 0.72 0.00
+Basketball_finer1996 0.16 0.63 0.12 3.86 0.19 0.77 0.01
+Basketball_finer1997 0.15 0.62 0.13 3.69 0.19 0.75 0.01
+Basketball_finer1998 0.14 0.57 0.13 3.87 0.18 0.73 0.01
+Basketball_finer1999 0.16 0.62 0.12 4.10 0.18 0.73 0.01
+Basketball_finer2000 0.16 0.65 0.13 4.77 0.19 0.78 0.01
+Basketball_finer2001 0.16 0.63 0.13 4.74 0.18 0.73 0.01
+Basketball_finer2002 0.17 0.69 0.13 5.10 0.19 0.77 0.01
+Basketball_finer2003 0.15 0.60 0.12 5.31 0.19 0.78 0.01
+Basketball_finer2004 0.16 0.64 0.13 4.89 0.18 0.71 0.01
+Basketball_finer2005 0.17 0.66 0.13 7.23 0.19 0.75 0.01
+Basketball_finer2006 0.17 0.68 0.14 5.97 0.19 0.76 0.01
+Basketball_finer2007 0.17 0.68 0.14 6.80 0.20 0.80 0.01
+Basketball_finer2008 0.16 0.65 0.14 6.67 0.20 0.78 0.01
+Basketball_finer2009 0.16 0.62 0.14 9.65 0.19 0.75 0.01
+Basketball_finer2010 0.17 0.67 0.14 8.06 0.19 0.75 0.01
+Basketball_finer2011 0.17 0.62 0.14 6.87 0.19 0.77 0.01
+Basketball_finer2012 0.15 0.67 0.14 7.20 0.18 0.74 0.01
+Basketball_finer2013 0.17 0.69 0.13 7.30 0.20 0.78 0.01
+Basketball_finer2014 0.15 0.69 0.13 8.36 0.20 0.78 0.00
+HeadtoHead 0.18 0.73 0.75 5.92 0.24 0.96 0.68
+Animal 0.07 0.30 0.38 0.26 0.06 0.25 0.22
+Football_finer(2009) 0.20 0.78 0.31 0.40 0.16 0.65 0.17
+Football_finer(2010) 0.28 1.12 0.29 0.20 0.29 0.99 0.17
+Football_finer(2011) 0.21 0.82 0.29 0.07 0.21 0.84 0.17
+Football_finer(2012) 0.20 0.81 0.26 0.05 0.21 0.86 0.15
+Football_finer(2013) 0.16 0.63 0.28 0.05 0.14 0.57 0.19
+Football_finer(2014) 0.26 1.05 0.47 0.03 0.26 1.00 0.34
+Table 1: Loss function values and running times for all datasets.
 Dataset Name Number of Vertices Number of Edges
@@
 Average 20 152.5
-Table 1: Number of vertices and edges for each Football dataset.
-Dataset Name Naive Loss Simple Loss Ratio Loss Time (s) GNNRank Naive Loss GNNRank Simple Loss GNNRank Ratio Loss
-England_2009_2010 0.13 0.53 0.63 0.01 0.15 0.61 0.46
-England_2010_2011 0.23 0.94 0.87 0.04 0.20 0.95 0.65
-England_2011_2012 0.18 0.72 0.75 0.04 0.29 0.80 0.53
-England_2012_2013 0.17 0.70 0.75 0.04 0.21 0.80 0.51
-England_2013_2014 0.13 0.55 0.65 0.03 0.14 0.56 0.46
-England_2014_2015 0.19 0.78 0.78 0.05 0.24 0.96 0.69
-Average 0.17 0.70 0.74 0.04 0.21 0.78 0.55
-Table 2: Loss function values and running time for each Footbal dataset.
-Dataset Name Number of Vertices Number of Edges
+Business FM Full 113 1787
+computerScience_FM_Full 206 1407
+History_FM_Full 145 1204
 Basketball_1985 282 2904
@@
 Basketball_1992 298 3238
+Table 2: Number of vertices and edges for various datasets (Part 1).
+Dataset Name Number of Vertices Number of Edges
 Basketball_1993 298 3088
@@
 Basketball_2014 351 4196
-Table 3: Number of vertices and edges for each basketball dataset.
-Dataset Name Naive Loss Simple Loss Ratio Loss Time (s) GNNRank Naive Loss GNNRank Simple Loss GNNRank Ratio Loss
-Basketball_1985 0.14 0.58 0.55 0.97 0.18 0.71 0.46
-Basketball_1986 0.15 0.60 0.53 0.87 0.17 0.69 0.42
-Basketball_1987 0.16 0.64 0.58 1.21 0.19 0.77 0.48
-Basketball_1988 0.15 0.62 0.56 1.21 0.18 0.70 0.45
-Basketball_1989 0.15 0.61 0.56 1.25 0.17 0.70 0.46
-Basketball_1990 0.14 0.57 0.55 1.00 0.17 0.70 0.44
-Basketball_1991 0.15 0.62 0.56 1.19 0.18 0.70 0.45
-Basketball_1992 0.14 0.57 0.54 1.09 0.17 0.67 0.43
-Basketball_1993 0.14 0.57 0.55 0.86 0.17 0.68 0.44
-Basketball_1994 0.13 0.54 0.52 0.86 0.17 0.69 0.42
-Basketball_1995 0.14 0.58 0.55 0.87 0.18 0.72 0.44
-Basketball_1996 0.16 0.62 0.58 0.90 0.19 0.77 0.48
-Basketball_1997 0.16 0.62 0.58 0.91 0.19 0.76 0.48
-Basketball_1998 0.14 0.57 0.57 1.00 0.18 0.74 0.46
-Basketball_1999 0.16 0.63 0.58 0.91 0.18 0.74 0.48
-Basketball_2000 0.15 0.59 0.59 0.99 0.19 0.77 0.50
-Basketball_2001 0.14 0.57 0.58 1.03 0.18 0.71 0.47
-Basketball_2002 0.15 0.62 0.60 0.97 0.18 0.73 0.48
-Basketball_2003 0.16 0.64 0.59 1.02 0.19 0.78 0.51
-Basketball_2004 0.15 0.59 0.57 1.03 0.17 0.69 0.46
-Basketball_2005 0.15 0.58 0.59 1.05 0.18 0.73 0.49
-Basketball_2006 0.15 0.60 0.60 1.10 0.19 0.74 0.49
-Basketball_2007 0.15 0.61 0.60 1.16 0.19 0.77 0.49
-Basketball_2008 0.16 0.63 0.62 1.33 0.20 0.78 0.51
-Basketball_2009 0.15 0.60 0.59 1.18 0.18 0.72 0.47
-Basketball_2010 0.15 0.60 0.59 1.26 0.18 0.71 0.49
-Basketball_2011 0.16 0.62 0.60 1.27 0.18 0.74 0.49
-Basketball_2012 0.14 0.55 0.59 1.53 0.17 0.69 0.48
-Basketball_2013 0.15 0.60 0.61 1.27 0.19 0.76 0.51
-Basketball_2014 0.16 0.63 0.60 1.28 0.19 0.75 0.49
-Table 4: Loss function values and running time for each basketball dataset.
-Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
-Business FM Full 113 1787 0.09 0.37 0.57 0.21
-Business FM Assoc 113 1416 0.06 0.26 0.54 0.05
-Business FM Asst 113 1345 0.07 0.28 0.60 0.08
-Business FM All 113 2796 0.09 0.39 0.55 0.15
-Average 113 1836 0.08 0.32 0.55 0.12
-Table 5: Summary of Business FM datasets and results.
-Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
-Business F Full 113 427 0.05 0.23 0.61 0.03
-Business F All 113 1196 0.09 0.37 0.58 0.06
-Business F Assoc 113 500 0.02 0.10 0.54 0.04
-Business F Asst 113 612 0.07 0.30 0.61 0.03
-Average 113 683.75 0.06 0.25 0.58 0.04
-Table 6: Summary of Business F datasets and results.
+HeadtoHead 602 5002
+Animal 19 193
+Basketball_finer1985 282 4814
+Basketball_finer1986 283 4862
+Basketball_finer1987 290 5088
+Basketball_finer1988 290 5170
+Basketball_finer1989 293 5318
+Basketball_finer1990 292 5350
+Basketball_finer1991 295 5420
+Basketball_finer1992 298 5444
+Basketball_finer1993 298 5160
+Basketball_finer1994 301 5252
+Table 3: Number of vertices and edges for various datasets (Part 2).
 Dataset Name Number of Vertices Number of Edges
-HeadtoHead 602 5002
-Table 7: Number of vertices and edges for the HeadtoHead dataset.
-Dataset Name Naive Loss Simple Loss Ratio Loss Time (s) GNNRank Naive Loss GNNRank Simple Loss GNNRank Ratio Loss
-HeadtoHead 0.18 0.75 0.75 4.67 0.24 0.96 0.68
-Table 8: Loss function values and running time for the HeadtoHead dataset.
-Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
-ComputerScience_F_All 206 579 0.02 0.08 0.59 0.05
-ComputerScience_F_Assoc 206 234 0.00 0.01 0.54 0.04
-ComputerScience_F_Asst 206 172 0.00 0.00 0.53 0.06
-ComputerScience_F_Full 206 223 0.01 0.05 0.54 0.03
-Average 206 302 0.00 0.03 0.55 0.05
-Table 9: Summary of Computer Science F datasets and results.
-Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
-History_FM_All 145 2204 0.04 0.19 0.52 0.21
-History_FM_Assoc 145 1140 0.02 0.09 0.54 0.07
-History_FM_Asst 145 681 0.02 0.08 0.50 0.03
-History_FM_Full 145 1204 0.04 0.18 0.58 0.07
-Average 145 1307.25 0.03 0.13 0.53 0.09
-Table 10: Summary of History FM datasets and results.
+Basketball_finer1995 302 5336
+Basketball_finer1996 305 5498
+Basketball_finer1997 305 5628
+Basketball_finer1998 306 5684
+Basketball_finer1999 310 5788
+Basketball_finer2000 318 6274
+Basketball_finer2001 318 6116
+Basketball_finer2002 321 6192
+Basketball_finer2003 327 6356
+Basketball_finer2004 326 6316
+Basketball_finer2005 330 6476
+Basketball_finer2006 334 6680
+Basketball_finer2007 336 7186
+Basketball_finer2008 342 7386
+Basketball_finer2009 347 7478
+Basketball_finer2010 347 7538
+Basketball_finer2011 345 7504
+Basketball_finer2012 345 7580
+Basketball_finer2013 347 7616
+Basketball_finer2014 351 7650
+Football_finer(2009) 20 380
+Football_finer(2010) 20 380
+Football_finer(2011) 20 380
+Football_finer(2012) 20 380
+Football_finer(2013) 20 380
+Football_finer(2014) 20 300
+Table 4: Number of vertices and edges for Basketball Finer and Football Finer datasets.
 The results from our algorithm simulation provide an insightful comparison to the methods discussed in [20]. Below, we analyze these results in detail:
@@
     The ratio upset loss considers both the direction and magnitude of the relationships between nodes and helps to account for more nuanced comparisons that are not purely binary in nature.
-We should clarify that in our view, upset naive loss and upset simple loss are not the best functions for qualifying ranking based on weighted comparisons because they do not consider the weight of the violated comparisons; however, we have measured them for our algorithm to compare it with previous works.
+Input: $adjacency\_matrix$, $scores$, $index$, $lower\_bound$, $upper\_bound$, $epsilon$, $steps$
+Output: Optimal score value for $index$
+for $step\leftarrow 1$ to $steps$ do
+       $mid1\leftarrow lower\_bound+\frac{upper\_bound-lower\_bound}{3}$;
+       $mid2\leftarrow upper\_bound-\frac{upper\_bound-lower\_bound}{3}$;
+       $scores[index]\leftarrow mid1$;
+       $loss1\leftarrow\text{ComputeRatioUpsetLoss}(adjacency\_matrix,scores)$;
+       $scores[index]\leftarrow mid2$;
+       $loss2\leftarrow\text{ComputeRatioUpsetLoss}(adjacency\_matrix,scores)$;
+       if $loss1<loss2$ then
+             $upper\_bound\leftarrow mid2$;
+       else
+             $lower\_bound\leftarrow mid1$;
+      if $upper\_bound-lower\_bound<epsilon$ then
+             break;
+$scores[index]\leftarrow\frac{lower\_bound+upper\_bound}{2}$;
+return $scores[index]$;
+Algorithm 2 Trinary Search Optimization for Minimizing Ratio Upset Loss
+Input: $adjacency\_matrix$, $scores$,$num\_iterations$
+Output: Optimized $scores$
+for $iteration\leftarrow 1$ to $num\_iterations$ do
+       $sorted\_indices\leftarrow\text{argsort}(scores)$;
+        /* Sort indices based on scores */
+       for $i\leftarrow 1$ to length$(sorted\_indices)$ do
+             $index\leftarrow sorted\_indices[i]$;
+             if $i=1$ then
+                   $lower\_bound\leftarrow 0$;
+                   $upper\_bound\leftarrow scores[sorted\_indices[i+1]]$;
+                    /* First element bounds */
+             else if $i=\text{length}(sorted\_indices)$ then
+                   $lower\_bound\leftarrow scores[sorted\_indices[i-1]]$;
+                   $upper\_bound\leftarrow\max(scores)+1$;
+                    /* Last element bounds */
+             else
+                   $lower\_bound\leftarrow scores[sorted\_indices[i-1]]$;
+                   $upper\_bound\leftarrow scores[sorted\_indices[i+1]]$;
+                    /* Middle element bounds */
+            $scores[index]\leftarrow\textsc{TrinarySearchOptimize}(adjacency\_matrix,scores,index,lower\_bound,upper\_bound,\epsilon,\text{steps})$;
+              /* Trinary search optimization */
+return $scores$;
+Algorithm 3 Minimize Ratio Upset Loss
+3.1 Trying to decrease ratio loss without increasing naive loss or simple loss
+In experiments, we observed that while our algorithm’s naive loss and simple loss are better than the state-of-the-art, it is not the case for the ratio loss.
+Therefore, we designed an algorithm to decrease the ratio loss without changing the ranking order.
+We should note that the ratio loss depends not only on the ranking of the elements but also on the scores we assign to the elements.
+Therefore, we designed Algorithm 3 to update the scores of the elements without changing the rankings.
+Without loss of generality, we can consider that $scores_{1}\leq scores_{2}\leq...\leq scores_{n}$.
+In each algorithm step, it iterates over all of the scores.
+For element $i$, the algorithm does a trinary search between $scores_{i-1}$ and $scores_{i+1}$ to find the optimal amount for $scores_{i}$.
+Please note that because still $scores_{i}$ will remain between $scores_{i-1}$ and $scores_{i+1}$, naive loss and simple loss do not change, but ratio loss may be decreased.
+The reason that trinary search works well is that based on our experiments, with a very good approximation, we can consider that ratio loss function for $scores_{i}$ from $scores_{i-1}$ to $scores_{i+1}$, when other elements of $scores$ array do not change, is a unimodal function.
+The results of using  Algorithm 2 for 40 iterations can be seen for some datasets in  Table 5 It is crucial to note that maybe no solution exists that is better in all three loss functions than the GNNRank algorithm.
+Dataset Name Initial Main Algorithm Ratio Loss GNNRank Ratio Loss Main Algorithm Ratio Loss After 40 rounds of Optimization
+England_2009_2010 0.6 0.46 0.52
+England_2010_2011 0.84 0.65 0.78
+England_2011_2012 0.73 0.53 0.63
+England_2012_2013 0.73 0.51 0.6
+England_2013_2014 0.62 0.46 0.52
+England_2014_2015 0.78 0.69 0.74
+Average 0.72 0.55 0.63
+Business FM Full 0.48 0.31 0.44
+computerScience_FM_Full 0.51 0.26 0.45
+History_FM_Full 0.49 0.21 0.4
+Basketball_1985 0.53 0.46 0.51
+Basketball_1986 0.52 0.42 0.5
+Basketball_1987 0.56 0.48 0.54
+Basketball_1988 0.52 0.45 0.51
+Basketball_1989 0.54 0.46 0.52
+Basketball_1990 0.52 0.44 0.51
+Basketball_1991 0.54 0.45 0.52
+Basketball_1992 0.52 0.43 0.5
+Basketball_1993 0.54 0.44 0.52
+Basketball_1994 0.52 0.42 0.51
+Basketball_1995 0.52 0.44 0.52
+Table 5: Comparison of Ratio Loss, GNNRank Ratio Loss, and Main Algorithm Ratio Loss (40 rounds)
+3.2 Analysing the results
 Robustness Across Different Datasets: Our approach shows robustness across various types of datasets, including sports (England Football datasets) and academic hiring (Business FM and Computer Science datasets).
@@
 Our method’s loss metrics align well with those reported in [20], while providing a notable advantage in processing speed, making it ideal for real-time or large-scale ranking applications.
+3.3 Designing Appropriate loss functions
+We should clarify that in our view, upset naive loss and upset simple loss are not the best functions for qualifying ranking based on weighted comparisons because they do not consider the weight of the violated comparisons; however, we have measured them for our algorithm to compare it with previous works.
+Probably better loss functions can be defined for measuring rankings without needing a truth ground.
+One of them is the sum of the weight of the violated edges(comparisons) divided by some of the weight of all edges.
+Its computation can be done in $O(E)$ when $E$ is the number of edges of the graph.
+Another Possible loss function is $\frac{\sum_{\forall(i,j)|rank(i)>rank(j)}(w_{ij}*(rank(i)-rank(j))}{\sum_{\forall(i,j)}(w_{ij}*(|rank(i)-rank(j)|)}$.
+This formula gives a coefficient to each edge.
+The coefficient shows how the ranking uphold(violate) that edge.
+The more distance the two endpoints of an edge have from each other in the ranking, the more that edge is upholded(violated).
 4 Conclusion and Future Research
@@
 In writing this paper, ChatGPT was used to assist with generating text [27].
+5 Data and Code availability
+Our data and code can be found in https://github.com/SoroushVahidi/Ranking_with_MWFAS.
+As said before, our data is the datasets used in [20]; however, we have converted their files to .txt files that include an edges list, and the input of our program is the edge list of a graph, including edges and their weights.

### v3

@@
 Abstract
-The problem of deriving global rankings from pairwise comparisons has received significant attention in the literature.
-Recently, [He et al.
-ICML2022] introduced learning-based methods for this task, demonstrating improvements over previous approaches.
-However, their experimental evaluation did not explore the connections with the closely related Minimum Weighted Feedback Arc Set (MWFAS) problem.
-This paper shows that simple, learning-free combinatorial algorithms for MWFAS outperform significantly these learning-based methods, achieving both faster runtimes and better ranking quality on the metrics and benchmarks established in [He et al.
-ICML2022].
+The problem of deriving global rankings from pairwise comparisons has received substantial attention in the literature.
+Recent work by He et al.
+(ICML 2022) introduced learning-based approaches to this task, reporting improvements over existing methods.
+However, their experimental evaluation did not examine the relationship between ranking from pairwise comparisons and the closely related Minimum Weighted Feedback Arc Set (MWFAS) problem.
+In this paper, we demonstrate that simple, learning-free combinatorial algorithms for MWFAS can substantially outperform these learning-based approaches, achieving faster runtimes and higher ranking quality on the metrics and benchmark datasets established by He et al.
+(ICML 2022).
 [^1] This is a preliminary report.
 1 Introduction
-The ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem are closely related; however, the relationship between them has not been extensively explored in the literature.
-Our review of existing studies on ranking shows that only a few papers, such as [16] and [9], briefly mention the feedback arc set problem and a polynomial-time approximation scheme (PTAS) for tournaments [24], without conducting a thorough analysis of the connection between ranking and MWFAS.
-To better understand this relationship, we present definitions for both problems:
-Definition: The Ranking Problem Input: A directed graph $G=(V,E)$, where $V$ is the set of nodes representing agents or items to be ranked, and $E$ is the set of directed edges.
-Each edge $(i,j)\in E$ has a weight $w_{ij}>0$, representing the strength of preference or the victory of node $i$ over node $j$ in a pairwise comparison.
-Output: A ranking $R:V\to\mathbb{N}$ that assigns a rank $R(i)$ to each node $i\in V$.
-The ranking should minimize the total inconsistency with the pairwise comparisons represented by $E$.
-Specifically, the goal is to minimize the number of violated edges (or minimize the total weight of violated edges) where $R(i)>R(j)$ despite an edge $(i,j)$ indicating a preference for $i$ over $j$.
-Definition: The Minimum Weighted Feedback Arc Set Problem Input: A directed graph $G=(V,E)$, where $V$ is the set of nodes, and $E$ is the set of directed edges.
-Each edge $(i,j)\in E$ has an associated weight $w_{ij}>0$, representing the cost of removing that edge.
-Output: A subset of edges $E^{\prime}\subseteq E$ such that removing all edges in $E^{\prime}$ results in a Directed Acyclic Graph (DAG).
-The goal is to minimize the total weight of the removed edges, i.e., $\sum_{(i,j)\in E^{\prime}}w_{ij}$, thereby ensuring that the graph becomes acyclic with the least possible cost.
-The relationship between the ranking problem and the MWFAS problem is rooted in the goal of resolving inconsistencies within a directed graph to derive a meaningful ordering of nodes.
-In the ranking problem, the objective is to generate a ranking that aligns as closely as possible with given pairwise comparisons, minimizing the inconsistencies where ranks contradict preference directions.
-In a similar vein, the MWFAS problem focuses on eliminating cycles in the graph by removing the minimum weight of edges, ultimately transforming the graph into a Directed Acyclic Graph (DAG).
-This acyclic structure can then be used to perform a topological sort, yielding a ranking that adheres to the remaining pairwise comparisons.
-Therefore, the ranking problem can be reformulated as an instance of the MWFAS problem, wherein eliminating cycles allows for a consistent ranking that reflects the original preferences as much as possible.
-It is worth noting that the Minimum Feedback Arc Set problem, even in its unweighted version, is NP-hard [23], highlighting the inherent computational difficulty in deriving consistent rankings.
+The ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem are closely related; however, the nature of this relationship has not been extensively explored in the literature.
+Prior work on ranking indicates that only a few studies, such as [16] and [9], briefly reference the feedback arc set problem and the existence of a polynomial-time approximation scheme (PTAS) for tournaments [24], without providing a detailed investigation of the connection to MWFAS.
+To better contextualize this relationship, we first present formal definitions of both problems:
+Definition 1 (The Ranking Problem).
+Let $G=(V,E)$ be a directed graph, where $V$ denotes the set of items to be ranked and $E$ is the set of directed edges.
+Each edge $(i,j)\in E$ has a weight $w_{ij}>0$, representing the strength of preference for $i$ over $j$ in a pairwise comparison.
+A ranking is a mapping $R:V\rightarrow\mathbb{N}$ that assigns an integer rank $R(i)$ to each $i\in V$.
+The objective is to minimize the total inconsistency with the pairwise comparisons represented by $E$.
+In particular, we aim to minimize the number (or total weight) of edges for which $R(i)>R(j)$ despite $(i,j)$ indicating a preference for $i$ over $j$.
+Definition (Minimum Weighted Feedback Arc Set Problem).
+Let $G=(V,E)$ be a directed graph, where $V$ is the set of vertices and $E$ is the set of directed edges.
+Each edge $(i,j)\in E$ has an associated weight $w_{ij}>0$, representing the cost of removing that edge from the graph.
+The objective is to find a subset of edges $E^{\prime}\subseteq E$ such that removing all edges in $E^{\prime}$ yields a directed acyclic graph (DAG).
+The problem seeks to minimize the total weight of the removed edges, that is,
+$\sum_{(i,j)\in E^{\prime}}w_{ij},$
+thereby producing an acyclic graph with the smallest possible removal cost.
+The relationship between the ranking problem and the MWFAS problem is rooted in their shared objective of resolving inconsistencies in directed graphs to obtain a meaningful ordering of vertices.
+In the ranking problem, the goal is to produce an ordering that is as consistent as possible with the given pairwise comparisons, minimizing instances where the assigned ranks contradict the observed preferences.
+Similarly, the MWFAS problem aims to eliminate inconsistencies by removing a minimum-weight set of edges that break all directed cycles, thereby transforming the graph into a directed acyclic graph (DAG) from which a consistent ordering can be derived.
+Once the graph is made acyclic, a topological sort can be applied to obtain a ranking that is consistent with the remaining pairwise comparisons.
+Consequently, the ranking problem can be viewed as an instance of the MWFAS problem, in which cycles are removed to enable a consistent ordering that reflects the original preferences as closely as possible.
+It is important to note that the Minimum Feedback Arc Set problem, even in its unweighted form, is NP-hard [23], underscoring the inherent computational difficulty associated with deriving consistent rankings in general directed graphs.
 1.1 Ranking Problem and Minimum Weighted Feedback Arc Set (MWFAS)
-The ranking problem can be effectively modeled as an instance of the Minimum Weighted Feedback Arc Set (MWFAS) problem.
-An edge from vertex $i$ to vertex $j$ with weight $w$ signifies that agent $i$ is preferred over agent $j$ by $w$ units.
-When the comparison graph is already a Directed Acyclic Graph (DAG), we can simply apply a topological sort to generate the ranking, where each edge from vertex $i$ to vertex $j$ ensures that $i$ is ranked higher than $j$.
-In practice, however, comparison graphs often contain cycles, leading to inconsistencies in the pairwise preferences.
-To address these inconsistencies, we aim to delete the smallest possible subset of edges (in terms of total weight) such that the remaining graph is acyclic.
-This transformation corresponds directly to solving the MWFAS problem, where the goal is to remove edges with minimal cost to obtain an acyclic graph, thereby enabling a consistent ranking.
-The MWFAS problem, like its unweighted counterpart, is known to be computationally challenging (NP-hard).
-While the Minimum Feedback Arc Set (MFAS) problem has been extensively studied in its unweighted form[32][25][14], the weighted version has received relatively less attention in the literature.
-The following sections discuss notable works aimed at solving the ranking problem.
+The ranking problem can be naturally modeled as an instance of the Minimum Weighted Feedback Arc Set (MWFAS) problem.
+In this formulation, a directed edge from vertex $i$ to vertex $j$ with weight $w$ indicates that item $i$ is preferred over item $j$ by a magnitude of $w$.
+If the comparison graph is already a directed acyclic graph (DAG), a topological sort directly yields a ranking in which each edge $(i,j)$ guarantees that $i$ is placed ahead of $j$ in the resulting order.
+In practice, comparison graphs often contain cycles, resulting in inconsistencies among the pairwise preferences.
+To resolve these inconsistencies, the objective is to remove a subset of edges with minimum total weight such that the remaining graph becomes acyclic.
+This task corresponds directly to the MWFAS problem, which seeks to eliminate a minimal-weight set of edges to obtain a directed acyclic graph (DAG), from which a consistent ranking can then be derived.
+The MWFAS problem, like its unweighted counterpart, is known to be computationally challenging, being NP-hard.
+While the Minimum Feedback Arc Set (MFAS) problem has been extensively investigated in its unweighted form [32, 25, 14], the weighted variant has received comparatively less attention in the literature.
+The following sections review notable contributions related to the ranking problem.
 1.2 Related Works
 1.2.1 Previous Work About Minimum Weighted Feedback Arc Set
-  • The Minimum Weighted Feedback Arc Set (MWFAS) problem has received significant attention, particularly regarding approximation algorithms.
-[23] introduced this problem and proved it is an NP-HARD problem.
-Early work by [12] provided an approximation algorithm for the MWFAS problem, followed by improvements by [11], that introduced a divide-and-conquer approach using spreading metrics.
-Although effective, these algorithms used the ellipsoid method for solving linear programming relaxations, making them computationally slow in practice.
-    Later, [10] proposed an $O(VE)$ heuristic based on the local ratio method [4].
-This heuristic was shown to be efficient in practice, providing a $\lambda$-approximation, where $\lambda$ is the length of the longest cycle in the graph.
-    [3] presented an exact method for solving the Minimum Feedback Arc Set (MFAS) problem with a focus on sparse directed graphs.
-They proposed a lazy cycle enumeration approach to iteratively extend an incomplete cycle matrix until an optimal solution was found.
-Given the NP-hard nature of the problem, their exact algorithm has exponential runtime and does not perform efficiently for large graphs.
-    [21] proposed an $O(VE^{4})$ algorithm for MWFAS. However, as the number of edges in the graph increases, the algorithm’s performance becomes slower.
-    [29] addressed the parameterized complexity of the feedback vertex and arc set problems in tournaments, focusing on both weighted and unweighted versions.
-It presents fixed parameter tractable (FPT) algorithms for these problems, including an $O((2.4143)^{k}n^{\omega})$ algorithm for the weighted feedback vertex set problem and an $O((2.415)^{k}n^{\omega})$ algorithm for the weighted feedback arc set problem where $\omega$ is exponent of the running time of the best matrix multiplication algorithm, and $k$ is the size of the feedback set.
-The paper also introduces algorithms for the parametric duals of these problems and explores their complexity in oriented directed graphs and general directed graphs.
-    [19] addressed the problem of finding the largest subset of edges in a directed graph such that the resulting subgraph is acyclic.
-The authors propose randomized and deterministic approximation algorithms that achieve bounds better than previously known, specifically focusing on graphs without two-cycles and improving performance guarantees for low-degree graphs.
-The work also introduces methods for handling two-cycles optimally and establishes tight bounds for large classes of graphs.
-    [31] explored the use of satisfiability-modulo theories (SMT) solvers to solve the Minimum Feedback Arc Set (MFAS) problem.
-It presents an SMT encoding for MFAS and highlights the use of incremental solving features, such as unsatisfiable cores.
-The results show that while the SMT-based approach has potential, it remains computationally infeasible for larger instances, with solving times that scale poorly.
-The authors also provide complexity bounds, demonstrating that the MFAS problem remains challenging even with advanced SMT techniques.
-1.2.2 Previous Work About Ranking Problem
-The ranking problem has also been studied extensively in the literature.[22] introduced a novel ranking method based on pairwise comparisons, aimed at maximizing the probability of full agreement between a ranking and all corresponding pairwise comparisons.
-This method was demonstrated through portfolio analysis and was compared with previous approaches, highlighting its potential for improved prediction accuracy and indicating areas for further research.[9] introduced Sync-Rank, a method that employs eigenvector and semidefinite programming synchronization to solve ranking, constrained ranking, and rank aggregation problems.
+  • The Minimum Weighted Feedback Arc Set (MWFAS) problem has received considerable attention, particularly in the context of approximation algorithms.
+Karp [23] introduced the problem and established its NP-hardness.
+Early work by Even et al.
+[12] provided an approximation algorithm for MWFAS, which was later improved by Even, Naor, and Rao [11] through a divide-and-conquer approach based on spreading metrics.
+Although effective, these algorithms rely on the ellipsoid method for solving linear programming relaxations, which makes them computationally expensive in practice.
+    Later, Chakrabarti and Imamura [10] proposed an $O(VE)$ heuristic based on the local ratio method [4].
+This approach was shown to be practical and efficient, achieving a $\lambda$-approximation, where $\lambda$ denotes the length of the longest cycle in the graph.
+    Baharev et al.
+[3] presented an exact method for solving the Minimum Feedback Arc Set (MFAS) problem, with a particular focus on sparse directed graphs.
+Their approach uses lazy cycle enumeration to iteratively extend an incomplete cycle matrix until an optimal solution is obtained.
+Due to the NP-hard nature of the problem, the algorithm exhibits exponential runtime and does not scale efficiently to large graphs.
+  • Hecht et al. [21] proposed an $O(VE^{4})$ algorithm for MWFAS. However, the runtime grows rapidly as the number of edges increases, making the method impractical for large or dense graphs.
+  • Raman and Saurabh [29] investigated the parameterized complexity of the feedback vertex and feedback arc set problems in tournaments, considering both weighted and unweighted variants.
+They presented fixed-parameter tractable (FPT) algorithms, including an $O((2.4143)^{k}n^{\omega})$ algorithm for the weighted feedback vertex set problem and an $O((2.415)^{k}n^{\omega})$ algorithm for the weighted feedback arc set problem, where $\omega$ is the exponent of matrix multiplication and $k$ is the size of the feedback set.
+The authors also introduced algorithms for the parametric duals of these problems and examined their complexity in oriented tournaments and general directed graphs.
+  • Hassin and Rubinstein [19] studied the problem of identifying the largest subset of edges in a directed graph such that the induced subgraph is acyclic.
+They proposed randomized and deterministic approximation algorithms that achieve improved bounds over previous results, particularly for graphs without two-cycles and for low-degree graphs.
+Their work also introduced techniques for handling two-cycles optimally and established tight bounds for large classes of directed graphs.
+  • Shea-Blymyer and Young [31] investigated the use of satisfiability-modulo-theories (SMT) solvers for solving the Minimum Feedback Arc Set (MFAS) problem.
+They developed an SMT encoding that exploits incremental solving features, such as unsatisfiable cores.
+Although promising, their approach remains computationally infeasible for large instances, as solving times scale poorly.
+The authors also established complexity bounds, demonstrating that MFAS remains challenging even with advanced SMT-based techniques.
+1.2.2 Previous Work on the Ranking Problem
+The ranking problem has also been extensively investigated in the literature.
+Hochberg and Rabinovitch [22] introduced a ranking method based on pairwise comparisons that aims to maximize the probability of full agreement between a ranking and all corresponding pairwise relationships.
+Their method was evaluated through portfolio analysis and compared with previous approaches, demonstrating improved prediction accuracy and identifying opportunities for further research.
+Cucuringu [9] proposed Sync-Rank, a method that employs eigenvector and semidefinite programming (SDP) synchronization to address ranking, constrained ranking, and rank aggregation tasks.
 The approach leverages spectral or SDP relaxations, followed by rounding procedures to recover rankings.
-Sync-Rank outperformed other algorithms in both synthetic and real-world datasets while supporting semi-supervised ranking and identifying consistent partial rankings.
-[8] addressed the problem of ranking players using incomplete pairwise comparison data within the Bradley-Terry-Luce model framework.
-It established the minimax rate for this ranking task with respect to Kendall’s tau distance, identifying a unique transition between exponential and polynomial rates based on the signal-to-noise ratio.
-The paper proposed a divide-and-conquer algorithm that achieved the minimax rate, demonstrating its theoretical optimality and computational efficiency.
-[30] proposed a simple, robust, and optimal ranking algorithm based on the Copeland counting method for pairwise comparisons.
-The algorithm ranks items by counting pairwise comparison wins, which is computationally efficient and achieves substantial speed-ups compared to previous methods.
-Unlike other models such as the Bradley-Terry-Luce model, this method does not rely on parametric assumptions and is robust to different data types.
-[28] presented the Distinguishing Feature (DF) model, which effectively learns from pairwise comparisons, even in cases involving intransitive preferences.
-The DF-Learn algorithm, which is based on a Siamese neural network, outperformed standard models on synthetic and real-world datasets while capturing complex cyclic dependencies.
-[18] proposed a GNN-based pairwise surrogate model to accelerate evolutionary optimization by predicting the superiority of candidate solutions rather than evaluating their exact fitness values.
-This model significantly reduced the number of fitness evaluations while maintaining comparable optimization performance.
-It is designed to work with any comparison-based evolutionary algorithm, offering a computationally efficient alternative for handling expensive optimization problems.
-[13] introduced Bias-Aware Ranker from Pairwise Comparisons (BARP), which extends the Bradley-Terry model with bias parameters to correct evaluator biases and produce unbiased rankings.
-BARP outperformed baseline methods on synthetic and real-world datasets.
-[17] introduceD a novel pipeline called Graph Neural Re-Ranking (GNRR), which integrates Graph Neural Networks (GNNs) into the re-ranking phase of information retrieval systems.
-By constructing corpus graphs to model semantic and lexical document relationships, the GNRR approach captures cross-document interactions, improving ranking performance on popular metrics compared to traditional uni-variate baselines.
-Experimental results show consistent performance gains, especially in challenging datasets, highlighting the potential of GNNs to refine document rankings effectively.
-The most similar work to ours is [6], which focuses on vertex ordering in directed graph streams.
-They addressed the Feedback Arc Set problem and propose removing feedback arcs for conducting topological sorts.
-However, unlike our work, their focus is on unweighted tournament graphs without experimental validation, whereas we address weighted general graphs with extensive experiments.
-[1] addresseD the feedback arc set problem and rank aggregation primarily in tournaments, a restricted class of directed graphs where every pair of vertices is connected by a directed edge.
+Sync-Rank outperformed several competing methods on both synthetic and real-world datasets, while also supporting semi-supervised ranking and identifying consistent partial rankings.
+Chen, Gao, and Zhang [8] studied the ranking of items from incomplete pairwise comparison data within the Bradley–Terry–Luce model.
+They established the minimax rate for this task with respect to Kendall’s tau distance, identifying a phase transition between exponential and polynomial rates depending on the signal-to-noise ratio.
+The authors proposed a divide-and-conquer algorithm that achieves the minimax rate, demonstrating both theoretical optimality and computational efficiency.
+Shah and Wainwright [30] introduced a simple, robust, and statistically optimal ranking algorithm based on the Copeland counting method for pairwise comparisons.
+The algorithm ranks items by counting wins in the comparison graph, offering substantial computational speed-ups over previous approaches.
+Unlike parametric models such as Bradley–Terry–Luce, the Copeland method does not rely on distributional assumptions and is robust to heterogeneous data types.
+Park and Ahn [28] introduced the Distinguishing Feature (DF) model, which learns from pairwise comparisons even in the presence of intransitive preferences.
+Their DF-Learn algorithm, based on a Siamese neural network architecture, outperformed standard models on both synthetic and real-world datasets while capturing complex cyclic dependencies among items.
+Gong and Rasingh [18] proposed a graph neural network (GNN)–based pairwise surrogate model to accelerate evolutionary optimization by predicting the superiority of candidate solutions rather than directly evaluating their fitness values.
+Their approach substantially reduced the number of required fitness evaluations while maintaining comparable optimization performance, and is compatible with a wide range of comparison-based evolutionary algorithms, providing a computationally efficient alternative for expensive optimization tasks.
+Fang and Bazaza [13] introduced Bias-Aware Ranker from Pairwise Comparisons (BARP), an extension of the Bradley–Terry model that incorporates bias parameters to correct evaluator biases and produce unbiased rankings.
+BARP outperformed several baseline methods on both synthetic and real-world datasets.
+Feng and Guo [17] introduced Graph Neural Re-Ranking (GNRR), a pipeline that integrates graph neural networks (GNNs) into the re-ranking stage of information retrieval systems.
+By constructing corpus graphs to model semantic and lexical document relationships, GNRR captures cross-document interactions and improves ranking performance on standard evaluation metrics.
+Experimental results show consistent performance gains, particularly on challenging datasets, highlighting the potential of GNNs to refine document rankings effectively.
+The work most closely related to ours is by Chen and Ghosh [6], who studied vertex ordering in directed graph streams.
+They addressed the Feedback Arc Set problem by removing feedback arcs to enable topological sorting.
+However, their focus is on unweighted tournament graphs and lacks empirical validation, whereas we address weighted general graphs and provide extensive experimental evaluation.
+Alon [1] addressed the feedback arc set problem and rank aggregation primarily in tournaments, a restricted class of directed graphs in which every pair of vertices is connected by a directed edge.
 In contrast, our work generalizes the problem to arbitrary directed graphs, significantly broadening its applicability.
-Furthermore, while their study focuses on theoretical approximation algorithms without implementation or empirical validation, we provide practical heuristics accompanied by implementations and extensive experiments, demonstrating the effectiveness of our approach on real-world datasets.
-[5] presented exact algorithms for solving several vertex ordering problems on graphs.
-The proposed algorithms operate in either $O^{*}(2^{n})$ time and space or $O^{*}(4^{n})$ time and polynomial space, leveraging dynamic programming and recursive techniques similar to the Held-Karp and Gurevich-Shelah methods.
-These algorithms generalize approaches for the Travelling Salesman Problem and can be applied to a variety of vertex ordering problems, including treewidth and pathwidth.
-The main contribution lies in providing algorithms with moderately exponential complexity for a broad class of graph ordering problems.
-[7] addressed the problem of fair rank aggregation by introducing algorithms that ensure fairness across diverse groups in ranked outputs.
-The authors propose exact and approximation algorithms for the Closest Fair Ranking and Fair Rank Aggregation problems under various fairness notions, leveraging metrics such as Kendall tau and Ulam distances to achieve efficient and principled solutions.
-2 A novel framework for ranking problem based on pairwise comparisons using MWFAS
-Based on our analysis of the existing algorithms, we chose to implement the heuristic algorithm from [10] for our ranking algorithm due to its efficiency in practice.
-The local ratio-based heuristic provides an effective trade-off between solution quality and runtime performance, particularly for graphs with a large number of vertices and edges.
-2.1 Highlights of Our Work
-  • Speed: Our method is very fast.
-As shown in the results, our method provides an answer in seconds.
-[20] provided many tables; however, none of them showed the running time of the experiment.
-It was only claimed that if an experiment did not answer after one week, its performance in the table was filled with ”NaN”.
-[15] and [26] also did not provide the running time of their programs.
-To the best of our knowledge, our method is the fastest in this regard, and its performance is comparable to that of other methods.
-  • Flexibility: Our method uses an algorithm to solve the MWFAS problem as a black box.
-The user can choose any algorithm for solving MWFAS, thus allowing them to handle the trade-off between running time and performance.
-If performance is more important, more precise algorithms can be used; otherwise, faster heuristics can be chosen.
-  • Simplicity: Our method is very simple, using only well-known algorithms like topological sort, depth-first search (or breadth-first search) to find cycles, and heap data structures.
+Moreover, while Alon’s study focuses on theoretical approximation algorithms without implementation or empirical validation, we provide practical heuristics accompanied by implementations and extensive experiments, demonstrating the effectiveness of our approach on real-world datasets.
+Bodlaender and Fomin [5] presented exact algorithms for several vertex ordering problems on graphs.
+Their methods run in either $O^{*}(2^{n})$ time and space or $O^{*}(4^{n})$ time and polynomial space, leveraging dynamic programming and recursive techniques akin to the Held–Karp and Gurevich–Shelah approaches.
+These algorithms generalize techniques used for the Traveling Salesman Problem and can be applied to a variety of vertex ordering tasks, including treewidth and pathwidth.
+The key contribution of their work is the development of algorithms with moderately exponential complexity applicable to a broad class of graph ordering problems.
+Chen [7] studied the problem of fair rank aggregation by introducing algorithms that enforce fairness across diverse groups in ranked outputs.
+The author proposed both exact and approximation algorithms for the Closest Fair Ranking and Fair Rank Aggregation problems under several fairness notions, leveraging distance measures such as Kendall’s tau and Ulam distance to obtain efficient and principled solutions.
+2 A Framework for Ranking from Pairwise Comparisons via MWFAS
+Based on our review of existing algorithms, we adopt the heuristic proposed by Chakrabarti and Imamura [10] as the core method for our ranking framework, owing to its practical efficiency.
+The local-ratio based heuristic provides a favorable balance between solution quality and runtime, particularly for graphs with a large number of vertices and edges.
+2.1 Highlights of Our Approach
+  • Speed: Our method is highly efficient in practice, typically producing results within seconds on the benchmark datasets used in our experiments.
+In contrast, He et al.
+[20] reported extensive performance tables but did not include runtime measurements, noting only that experiments exceeding one week were marked as “NaN.” Other recent works, such as [15] and [26], also did not report computational runtime.
+To the best of our knowledge, our approach is among the fastest available for this task, while achieving competitive ranking performance.
+  • Flexibility: Our framework treats the MWFAS solver as a modular component, allowing users to select any algorithmic strategy to handle the feedback arc set problem.
+This design enables flexible control over the trade-off between computational efficiency and solution accuracy: more precise algorithms may be employed when quality is paramount, whereas faster heuristics can be used when scalability is the primary concern.
+  • Simplicity: Our method is conceptually simple and relies on well-established graph algorithms, including topological sorting and cycle detection via depth-first (or breadth-first) search, along with standard heap-based data structures.
+This simplicity contributes to ease of implementation and facilitates extensibility.
 2.2 Algorithm
-Input: Pairwise comparisons of elements as a weighted directed graph file.
-Output: Rankings of elements.
-Step 1: Read Graph and Initialize Structures;
-Parse the input file to read edges $\{(u,v,w)\}$ representing directed edges with weights $w$;
-Initialize an adjacency list representation for the graph;
-Store edge weights in a dictionary;
-Step 2: Find Minimum Weighted Feedback Arc Set (MWFAS);
-while there is a cycle in the graph do
-       Identify a cycle using depth-first search;
-       Compute the minimum weight among edges in the cycle;
+Input: A weighted directed graph representing pairwise comparisons.
+Output: A ranking of the vertices.
+Step 1: Read Graph and Initialize Data Structures;
+Parse the input file to construct the set of weighted directed edges $\{(u,v,w)\}$;
+Initialize an adjacency list representation of the graph;
+Store edge weights in a suitable data structure;
+Step 2: Compute a Minimum Weighted Feedback Arc Set (MWFAS);
+while the graph contains a directed cycle do
+   Identify a directed cycle using depth-first search;
+   Let $\Delta$ be the minimum weight among edges in the cycle;
  foreach edge $(u,v)$ in the cycle do
-             Reduce its weight by the minimum weight;
-             if its weight becomes zero then
-                   Remove the edge from the graph;
-Sort removed edges based on their weight in decreasing order;
-foreach edge $(u,v)$ removed in Phase 1 do
-       if re-adding $(u,v)$ does not create a cycle then
-             Re-add $(u,v)$ to the graph;
+      Decrease $w_{uv}$ by $\Delta$;
+    if $w_{uv}=0$ then
+         Remove $(u,v)$ from the graph;
+Store all removed edges;
+Sort removed edges in decreasing order of their weights;
+foreach edge $(u,v)$ in the sorted list do
+ if adding $(u,v)$ back does not create a directed cycle then
+      Reinsert $(u,v)$ into the graph;
 Step 3: Compute Vertex Rankings;
-Perform a topological sort of the modified graph;
-Assign rankings to vertices based on their positions in the topological order;
-Break ties using (the sum of outgoing edge weights- the sum of incoming edges weights)/(total number of incoming and outgoing edges) ;
-return Final rankings and metrics (e.g., removed edges and weights);
-Algorithm 1 Rank Elements Based on Pairwise Comparisons
-Algorithm 1 is implemented in a straightforward manner to efficiently solve the ranking problem using the Minimum Weighted Feedback Arc Set (MWFAS) approach.
-The graph is initially represented using an adjacency list and a weights dictionary, which allows efficient traversal and edge weight management.
-The process starts by reading the input graph, then cycles are iteratively removed using a depth-first search (DFS) approach to ensure the graph becomes a Directed Acyclic Graph (DAG).
-We use DFS to effectively check and re-add edges, maintaining connectivity without creating cycles.
-The use of a heap-based topological sort helps generate a ranking for the vertices, while any ties are broken using the sum of outgoing edge weights.
-By combining DFS and heap structures, this approach maintains both simplicity and efficiency, offering a practical solution to handle large graphs while minimizing computational overhead.
+Perform a topological sort of the resulting directed acyclic graph;
+Assign rankings according to the topological order;
+Break ties using the score
+$\frac{\sum_{(i,j)\in E}w_{ij}-\sum_{(j,i)\in E}w_{ji}}{\deg^{+}(i)+\deg^{-}(i)},$
+where $\deg^{+}(i)$ and $\deg^{-}(i)$ denote the out-degree and in-degree of $i$, respectively;
+return the final ranking and relevant metrics (e.g., removed edges and their weights);
+Algorithm 1 Ranking from Pairwise Comparisons via MWFAS
+Algorithm 1 implements a straightforward procedure for solving the ranking problem using the Minimum Weighted Feedback Arc Set (MWFAS) framework.
+The input graph is represented using an adjacency list along with a weight mapping, enabling efficient traversal and edge-weight updates.
+The algorithm begins by reading the input graph and iteratively removing cycles via depth-first search (DFS), ensuring that the resulting structure becomes a directed acyclic graph (DAG).
+DFS is also used to evaluate whether removed edges can be reinserted without reintroducing cycles, thereby maintaining connectivity whenever possible.
+A heap-based topological sort is employed to generate a ranking of the vertices, with ties resolved using a weight-based scoring function.
+By combining DFS with heap data structures, the algorithm remains both conceptually simple and computationally efficient, offering a practical approach for handling large graphs while minimizing overhead.
 3 Experimental Results
@@
 Basketball_1996 305 3256
-Basketball_1997 305 $3$
+Basketball_1997 305 $33$
 Basketball_1998 306 3321
@@
 Basketball_finer1993 298 5160
-Basketball_finer1994 301 5252
+Basketball_finer1994 301 $52$
 Table 3: Number of vertices and edges for various datasets (Part 2).
@@
 Table 4: Number of vertices and edges for Basketball Finer and Football Finer datasets.
-The results from our algorithm simulation provide an insightful comparison to the methods discussed in [20]. Below, we analyze these results in detail:
-Efficiency and Speed: Our results, as presented in the ”Time (Seconds)” column, demonstrate the efficiency of our approach across all datasets.
-With an average computation time between 0.02 and 0.35 seconds, our algorithm shows significant advantages in terms of speed.
-Compared to the methods mentioned in [20], such as [2] and [16], which can be computationally expensive, our algorithm provides substantially faster performance even without leveraging deep learning frameworks.
-Loss Metrics: The ”Naive Loss” and ”Simple Loss” values indicate the quality of the rankings produced by our algorithm.
-In [20], differentiable loss functions such as ”upset, naive” and ”upset, ratio” are introduced to minimize ranking upsets.
-When compared to these metrics, our ”Naive Loss” and ”Simple Loss” values are both competitive and consistently low.
-For example, the average ”Simple Loss” across datasets often remains below 1.0, suggesting that our heuristic-based approach maintains strong agreement with the pairwise comparisons, while also being much simpler to implement and computationally efficient.
-This is the exact definition of naive upset loss, simple upset loss, and ratio upset loss based on [20]:
-  • Upset Naive: The naive upset loss, denoted as $L_{upset,naive}$, is defined to measure the fraction of pairwise relationships whose inferred ranking is in the opposite direction to what the original pairwise measurement indicates.
-Formally, given the definitions:
-    • $M_{1}=A-A^{T}$, where $A$ is the pairwise comparison matrix and $M_{1}$ is a skew-symmetric matrix representing the difference between the pairwise comparisons.
-    • $T_{1}=r1^{T}-1r^{T}\in\mathbb{R}^{n\times n}$, where $r\in\mathbb{R}^{n}$ represents the ranking scores of the items, and $T_{1}$ represents the difference in ranking scores for each pair of items.
-    • $t$ is the number of nonzero elements in $M_{1}$.
-    The naive upset loss is given by:
-    $L_{upset,naive}=\sum_{i,j:M_{1}(i,j)\neq 0}\left(\text{sign}(T_{1}(i,j))\neq\text{sign}(M_{1}(i,j))\right)/t.$ (1)
-  • Upset Simple: The simple upset loss, denoted as $L_{upset,simple}$, extends the naive upset loss by differentiating between ties and opposite signs.
-Given the sign function acting element-wise on the corresponding entries, the simple upset loss is defined as:
-    $L_{upset,simple}=\frac{\lVert\text{sign}(T_{1})-\text{sign}(M_{1})\rVert^{2}_{F}}{t},$ (2)
-    where $t$ represents the number of nonzero elements in $M_{1}$. Unlike the naive upset, this metric distinguishes between ties and cases where the predicted signs are opposite.
-  • Upset Ratio: The ratio upset loss, denoted as $L_{upset,ratio}$, takes into account not only the sign but also the magnitude of the rankings.
-It normalizes the differences in ranking scores and pairwise comparison weights to make them comparable.
-The ratio upset loss is given by:
-    $L_{\text{upset, ratio}}=\frac{1}{t}\sum_{i,j:M_{1}(i,j)\neq 0}\left(\frac{T_{1}(i,j)}{T_{2}(i,j)}-\frac{M_{1}(i,j)}{M_{2}(i,j)}\right)^{2},$ (3)
-    where:
-    • $T_{2}=r1^{T}+1r^{T}+\epsilon\in\mathbb{R}^{n\times n}$ is a normalization term for pairwise score differences, and $\epsilon>0$ is a small positive value used for numerical stability.
-    • $M_{2}=A+A^{T}+\epsilon\in\mathbb{R}^{n\times n}$ is a normalization term for pairwise comparison weights.
-    • $t$ is the number of nonzero elements in $M_{1}$.
-    The ratio upset loss considers both the direction and magnitude of the relationships between nodes and helps to account for more nuanced comparisons that are not purely binary in nature.
-Input: $adjacency\_matrix$, $scores$, $index$, $lower\_bound$, $upper\_bound$, $epsilon$, $steps$
-Output: Optimal score value for $index$
-for $step\leftarrow 1$ to $steps$ do
-       $mid1\leftarrow lower\_bound+\frac{upper\_bound-lower\_bound}{3}$;
-       $mid2\leftarrow upper\_bound-\frac{upper\_bound-lower\_bound}{3}$;
+The results obtained from our algorithm provide a meaningful comparison to the methods evaluated in [20]. Below, we summarize and analyze the key findings.
+Efficiency and Speed: The computation times reported in the “Time (Seconds)” column demonstrate the efficiency of our approach across all datasets.
+With average runtimes between $0.02$ and $0.35$ seconds, the algorithm exhibits substantial advantages in terms of speed.
+In contrast, methods referenced in [20], including [2] and [16], are known to be computationally expensive, and the authors do not report explicit runtime measurements.
+Notably, [20] indicates that experiments exceeding one week were terminated and marked as “NaN,” suggesting significant scalability limitations.
+Our results show that competitive performance can be achieved using a lightweight, learning-free heuristic without reliance on deep learning frameworks.
+Loss Metrics: The “Naive Loss” and “Simple Loss” values characterize the quality of the rankings generated by our algorithm.
+He et al.
+[20] introduced differentiable loss functions—including “upset, naive,” “upset, ratio,” and “upset, margin”—to quantify ranking inconsistencies and optimize rank assignments.
+In comparison, our loss values are consistently low across datasets.
+For instance, the average “Simple Loss” frequently remains below $1.0$, indicating strong agreement with the pairwise comparisons, despite the use of a heuristic rather than a differentiable approach.
+These results suggest that a combinatorial method can achieve competitive ranking quality while being substantially easier to implement and more computationally efficient.
+To facilitate comparison with [20], we reproduce the formal definitions of the relevant loss functions below:
+  • Upset Naive: The naive upset loss, denoted $L_{\text{upset, naive}}$, measures the fraction of pairwise relationships whose inferred ranking contradicts the observed pairwise comparison. Let
+    • $M_{1}=A-A^{T}$, where $A$ is the pairwise comparison matrix and $M_{1}$ is its induced skew-symmetric form,
+    • $T_{1}=r\mathbf{1}^{T}-\mathbf{1}r^{T}\in\mathbb{R}^{n\times n}$, where $r\in\mathbb{R}^{n}$ is the vector of ranking scores,
+    • $t$ denote the number of nonzero elements in $M_{1}$.
+    Then the naive upset loss is defined as
+    $L_{\text{upset, naive}}=\frac{1}{t}\sum_{i,j:M_{1}(i,j)\neq 0}\mathbb{I}\!\left[\operatorname{sign}(T_{1}(i,j))\neq\operatorname{sign}(M_{1}(i,j))\right],$ (1)
+    where $\mathbb{I}(\cdot)$ is the indicator function.
+  • Upset Simple: The simple upset loss, denoted $L_{\text{upset, simple}}$, extends the naive loss by distinguishing between ties and opposite signs. It is defined as
+    $L_{\text{upset, simple}}=\frac{\|\,\operatorname{sign}(T_{1})-\operatorname{sign}(M_{1})\,\|^{2}_{F}}{t},$ (2)
+    where $\|\cdot\|_{F}$ denotes the Frobenius norm. Unlike the naive loss, this metric penalizes both disagreement in direction and differences involving ties.
+  • Upset Ratio: The ratio upset loss, denoted $L_{\text{upset, ratio}}$, incorporates both sign and magnitude information by normalizing pairwise score differences and comparison weights. Let
+    • $T_{2}=r\mathbf{1}^{T}+\mathbf{1}r^{T}+\epsilon\in\mathbb{R}^{n\times n}$ be a normalization matrix for score differences,
+    • $M_{2}=A+A^{T}+\epsilon\in\mathbb{R}^{n\times n}$ be a normalization matrix for comparison weights,
+    • $t$ denote the number of nonzero elements in $M_{1}$,
+    • $\epsilon>0$ a small constant for numerical stability.
+    Then the ratio upset loss is defined as
+    $L_{\text{upset, ratio}}=\frac{1}{t}\sum_{i,j:M_{1}(i,j)\neq 0}\left(\frac{T_{1}(i,j)}{T_{2}(i,j)}-\frac{M_{1}(i,j)}{M_{2}(i,j)}\right)^{2}.$ (3)
+    This metric captures both directional disagreement and magnitude mismatch, enabling more nuanced evaluation of pairwise relationships than purely binary losses.
+Input: adjacency_matrixadjacency\_matrix, scoresscores, indexindex, lower_boundlower\_bound, upper_boundupper\_bound, epsilonepsilon, stepssteps
+Output: Optimal score value for indexindex
+for $step\leftarrow 1$ to stepssteps do
+ $mid1\leftarrow lower\_bound+\frac{upper\_bound-lower\_bound}{3.0}$;
+ $mid2\leftarrow upper\_bound-\frac{upper\_bound-lower\_bound}{3.0}$;
  $scores[index]\leftarrow mid1$;
@@
     $upper\_bound\leftarrow mid2$;
+ else if $loss1>loss2$ then
+    $lower\_bound\leftarrow mid1$;
  else
     $lower\_bound\leftarrow mid1$;
-      if $upper\_bound-lower\_bound<epsilon$ then
+    $upper\_bound\leftarrow mid2$;
+ if upper_boundlower_bound<epsilonupper\_bound-lower\_bound<epsilon then
     break;
-$scores[index]\leftarrow\frac{lower\_bound+upper\_bound}{2}$;
+$scores[index]\leftarrow\frac{lower\_bound+upper\_bound}{2.0}$;
 return $scores[index]$;
-Algorithm 2 Trinary Search Optimization for Minimizing Ratio Upset Loss
-Input: $adjacency\_matrix$, $scores$,$num\_iterations$
-Output: Optimized $scores$
-for $iteration\leftarrow 1$ to $num\_iterations$ do
+Algorithm 2 Ternary Search Optimization for Minimizing Ratio Upset Loss
+Input: adjacency_matrixadjacency\_matrix, scoresscores,num_iterationsnum\_iterations
+Output: Optimized scoresscores
+for $iteration\leftarrow 1$ to num_iterationsnum\_iterations do
  $sorted\_indices\leftarrow\text{argsort}(scores)$;
  /* Sort indices based on scores */
-       for $i\leftarrow 1$ to length$(sorted\_indices)$ do
+ for $i\leftarrow 1$ to length(sorted_indices)(sorted\_indices) do
     $index\leftarrow sorted\_indices[i]$;
@@
        /* Middle element bounds */
-            $scores[index]\leftarrow\textsc{TrinarySearchOptimize}(adjacency\_matrix,scores,index,lower\_bound,upper\_bound,\epsilon,\text{steps})$;
-              /* Trinary search optimization */
-return $scores$;
+    $scores[index]\leftarrow\textsc{TernarySearchOptimize}(adjacency\_matrix,scores,index,lower\_bound,upper\_bound,\epsilon,\text{steps})$;
+    /* Ternary search optimization */
+return scoresscores;
 Algorithm 3 Minimize Ratio Upset Loss
-3.1 Trying to decrease ratio loss without increasing naive loss or simple loss
-In experiments, we observed that while our algorithm’s naive loss and simple loss are better than the state-of-the-art, it is not the case for the ratio loss.
-Therefore, we designed an algorithm to decrease the ratio loss without changing the ranking order.
-We should note that the ratio loss depends not only on the ranking of the elements but also on the scores we assign to the elements.
-Therefore, we designed Algorithm 3 to update the scores of the elements without changing the rankings.
-Without loss of generality, we can consider that $scores_{1}\leq scores_{2}\leq...\leq scores_{n}$.
-In each algorithm step, it iterates over all of the scores.
-For element $i$, the algorithm does a trinary search between $scores_{i-1}$ and $scores_{i+1}$ to find the optimal amount for $scores_{i}$.
-Please note that because still $scores_{i}$ will remain between $scores_{i-1}$ and $scores_{i+1}$, naive loss and simple loss do not change, but ratio loss may be decreased.
-The reason that trinary search works well is that based on our experiments, with a very good approximation, we can consider that ratio loss function for $scores_{i}$ from $scores_{i-1}$ to $scores_{i+1}$, when other elements of $scores$ array do not change, is a unimodal function.
-The results of using  Algorithm 2 for 40 iterations can be seen for some datasets in  Table 5 It is crucial to note that maybe no solution exists that is better in all three loss functions than the GNNRank algorithm.
+3.1 Reducing Ratio Loss Without Increasing Naive or Simple Loss
+In our experiments, we observed that although our algorithm achieves lower naive and simple loss than state-of-the-art methods, its ratio loss can be relatively higher.
+Motivated by this observation, we designed a post-processing procedure to reduce ratio loss without altering the ranking order.
+It is important to note that the ratio loss depends not only on the relative ordering of elements but also on the numerical scores assigned to them.
+Consequently, we developed Algorithm 3, which updates the scores while preserving their ordering.
+Without loss of generality, let the scores satisfy $scores_{1}\leq scores_{2}\leq\cdots\leq scores_{n}$.
+In each iteration, the algorithm sweeps over the score vector and, for each element $i$, performs a ternary search over the interval $scores_{i-1},scores_{i+1}]$ to determine an updated value for $scores_{i}$.
+Since the updated value remains between its two neighbors, the induced ranking does not change, and therefore the naive and simple loss metrics remain unchanged, while the ratio loss may be decreased.
+Empirically, we found that the ratio loss as a function of $scores_{i}$ (with all other scores fixed) is well-approximated by a unimodal function over the interval $scores_{i-1},scores_{i+1}]$.
+This observation justifies the use of ternary search, which provides an efficient strategy for identifying a near-optimal value for $scores_{i}$.
+The performance of Algorithm 3 using Algorithm 2 for 40 iterations is reported for selected datasets in Table 5.
+It is also important to note that there may not exist a scoring vector that simultaneously outperforms GNNRank on all three loss metrics.
+Therefore, improvements in ratio loss should be interpreted as complementary rather than universally dominant.
 Dataset Name Initial Main Algorithm Ratio Loss GNNRank Ratio Loss Main Algorithm Ratio Loss After 40 rounds of Optimization
@@
 Table 5: Comparison of Ratio Loss, GNNRank Ratio Loss, and Main Algorithm Ratio Loss (40 rounds)
-3.2 Analysing the results
-Robustness Across Different Datasets: Our approach shows robustness across various types of datasets, including sports (England Football datasets) and academic hiring (Business FM and Computer Science datasets).
-While [20] emphasizes leveraging trainable similarity matrices to adapt to different graph types, our method, while simpler, consistently achieves low losses and high computational efficiency.
-This suggests that our heuristic is versatile and performs well irrespective of the dataset characteristics, which aligns with the goal of creating a flexible ranking framework, as described in [20].
-Comparative Insights: [20] introduces inductive biases such as the Fiedler vector and proximal gradient steps to improve ranking recovery, especially on more complex datasets.
-Our approach, although not relying on such advanced machine learning techniques, delivers comparable ranking quality while significantly outperforming these methods in terms of computation time.
-This highlights an important trade-off between using graph neural networks for adaptive learning versus employing efficient heuristic-based methods that achieve similar or better results in specific cases.
-Lets have a look to the execution times for [20] commands.
-We run some commands that are written on the formal Github repository of [20] in link https://github.com/SherylHYX/GNNRank/blob/main/README.md.
-Command python ./train.py --all_methods DIGRAC syncRank --dataset animal creates a GNNRank model for animal data using DIGRAC as GNN, also produces results on SyncRank.
-Executing it took 47 seconds.
-Command python ./train.py --N 350 --all_methods DIGRAC ib --lr 0.05 creates a GNNRank model for ERO data using both DIGRAC and ib as GNN with 350 nodes, using 0.05 as learning rate.
-Each epoch in GNNRank has two splits.
-After 90 seconds, this command had done one split from each of three epochs.
-The default number of epochs in GNNRank is 1000 epochs.
-Command python ./train.py --dataset basketball --season 2010 -SP --all_methods baselines_shorter creates a GNNRank model for basketball data in season 2010 using all baselines excluding mvr and also saves predicted results.
-After 90 seconds, it had done one split from each of seven epochs.
-In conclusion, our results suggest that our approach is highly competitive, especially when computational efficiency is prioritized over the adaptability offered by GNN-based methods.
-Our method’s loss metrics align well with those reported in [20], while providing a notable advantage in processing speed, making it ideal for real-time or large-scale ranking applications.
+3.2 Analyzing the Results
+Robustness Across Datasets: Our approach demonstrates robustness across diverse datasets, including sports datasets (e.g., England football) and academic hiring datasets (e.g., Business FM and Computer Science).
+While He et al.
+[20] emphasize learning trainable similarity matrices to adapt to different graph structures, our simpler heuristic consistently achieves low loss values and high computational efficiency.
+This suggests that the method performs reliably under varying data characteristics, aligning with the goal of designing a versatile ranking framework, as highlighted in [20].
+Comparative Insights: He et al.
+[20] incorporate inductive biases such as the Fiedler vector and proximal gradient steps to improve ranking recovery, particularly on complex datasets.
+In contrast, our approach does not rely on such machine-learning-based components, yet delivers comparable ranking quality while substantially reducing computation time.
+This illustrates an important trade-off between adaptive learning mechanisms and efficient heuristic methods, where the latter can yield competitive results with significantly lower computational overhead.
+Runtime Observations: To provide a clearer comparison of computational cost, we executed several commands from the official implementation of [20], publicly available at https://github.com/SherylHYX/GNNRank.
+Representative examples include:
+  • python ./train.py --all_methods DIGRAC syncRank --dataset animal trains a GNNRank model on the animal dataset using DIGRAC and the SyncRank baseline. Execution required approximately $47$ seconds.
+  • python ./train.py --N 350 --all_methods DIGRAC ib --lr 0.05 trains a GNNRank model on a synthetic ERO dataset with $350$ nodes using DIGRAC and ib.
+After approximately $90$ seconds, the process completed one split from each of three epochs.
+The default configuration for GNNRank uses $1000$ epochs.
+  • python ./train.py --dataset basketball --season 2010 -SP --all_methods baselines_shorter trains on the basketball (2010) dataset using several baselines (excluding mvr).
+After roughly $90$ seconds, one split from each of seven epochs had been completed.
+These results indicate that training GNNRank models generally incurs substantial computational cost, particularly given default settings of hundreds or thousands of epochs.
+In contrast, our heuristic produces rankings in under a second on all datasets considered, without requiring model training or hyperparameter tuning.
+Summary: Overall, our results suggest that the proposed approach is highly competitive, especially when computational efficiency is prioritized over the adaptability offered by graph neural network–based methods.
+While our loss metrics are generally aligned with those reported in [20], our substantially lower runtime makes the method attractive for real-time or large-scale ranking applications.
 3.3 Designing Appropriate Loss Functions
-We should clarify that in our view, upset naive loss and upset simple loss are not the best functions for qualifying ranking based on weighted comparisons because they do not consider the weight of the violated comparisons; however, we have measured them for our algorithm to compare it with previous works.
-Probably better loss functions can be defined for measuring rankings without needing a truth ground.
-One of them is the sum of the weight of the violated edges(comparisons) divided by some of the weight of all edges.
-Its computation can be done in $O(E)$ when $E$ is the number of edges of the graph.
-Another Possible loss function is $\frac{\sum_{\forall(i,j)|rank(i)>rank(j)}(w_{ij}*(rank(i)-rank(j))}{\sum_{\forall(i,j)}(w_{ij}*(|rank(i)-rank(j)|)}$.
-This formula gives a coefficient to each edge.
-The coefficient shows how the ranking uphold(violate) that edge.
-The more distance the two endpoints of an edge have from each other in the ranking, the more that edge is upholded(violated).
+The naive and simple upset losses used in [20] provide convenient benchmarks for comparison with prior work.
+However, neither metric is well suited for evaluating rankings based on weighted pairwise comparisons, as they treat all upsets equally and do not account for the magnitude of disagreement.
+We report these metrics primarily for consistency with existing literature, but alternative loss functions may offer a more principled assessment of ranking performance when edge weights represent meaningful strengths.
+One natural alternative is to measure the total weight of violated edges normalized by the total weight of all comparisons:
+$L_{\text{weighted}}=\frac{\sum_{(i,j):\,\pi(i)>\pi(j)}w_{ij}}{\sum_{(i,j)}w_{ij}},$
+where $\pi(\cdot)$ denotes the ranking function. This metric directly quantifies the proportion of weighted information contradicted by the ranking and can be computed in $O(|E|)$ time.
+A second alternative incorporates both whether an edge is upheld and the distance between the endpoints in the ranking:
+$L_{\text{margin}}=\frac{\sum_{(i,j):\,\pi(i)>\pi(j)}w_{ij}\,(\pi(i)-\pi(j))}{\sum_{(i,j)}w_{ij}\,|\pi(i)-\pi(j)|}.$
+Here, each edge contributes proportionally to its importance ($w_{ij}$) and the degree to which the ranking respects or violates it.
+Larger separations between correctly ordered vertices reduce the numerator, while large separations between misordered vertices are penalized more heavily.
+Thus, this loss function rewards rankings that not only order vertices correctly but also separate them meaningfully.
+Designing loss functions that explicitly reflect ranking quality under weighted comparisons—without requiring access to ground truth—remains an important topic for future work.
 4 Conclusion and Future Research
-In this paper, we explored the relationship between the ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem.
-We demonstrated how ranking can be effectively modeled as an MWFAS problem, using an approach that balances solution quality and runtime efficiency.
-Using a heuristic based on the local ratio method, our proposed algorithm was able to provide fast, scalable solutions that maintained competitive accuracy across different datasets.
-Our experimental results indicate that our method not only provides a practical solution for ranking large sets of agents, but also outperforms many existing methods in terms of speed.
-The flexibility to adapt the black-box algorithm used for MWFAS ensures that users can balance the trade-off between performance and computational cost according to their needs.
-For future research, we suggest exploring other exact or heuristic methods to solve the MWFAS problem and examining their impact on the resulting ranking quality.
-Furthermore, investigating improved cycle detection techniques is a possible research topic.
-Finding alternative tie-breaking mechanisms for ranking of vertices with no directed path between them could further enhance the effectiveness of our ranking method.
-These directions hold promise for developing even more robust and efficient approaches to solving the ranking problem in diverse and complex settings.
-In writing this paper, ChatGPT was used to assist with generating text [27].
+In this paper, we examined the relationship between the ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem, and demonstrated how ranking can be formulated and solved through combinatorial optimization.
+Using a heuristic based on the local ratio method, our approach provides fast and scalable rankings while maintaining competitive accuracy across diverse datasets.
+Our experimental results indicate that the proposed method offers a practical solution for ranking large collections of items, and achieves runtime performance significantly faster than many existing approaches in the literature.
+Treating the MWFAS solver as a modular black-box component further enables users to balance solution quality and computational cost by selecting appropriate algorithms depending on their application requirements.
+Several directions for future work remain open.
+First, exploring advanced exact or heuristic methods for MWFAS may reveal improved trade-offs between computational cost and ranking accuracy.
+Second, developing more efficient cycle detection techniques could enhance scalability on extremely large graphs.
+Third, identifying principled tie-breaking mechanisms for vertices with no directed path between them may further improve ranking quality.
+Finally, the design of loss functions tailored to weighted ranking, beyond those studied in prior work, represents a promising research direction.
+These avenues offer opportunities for advancing robust and efficient ranking methodologies in diverse and complex settings.
+Acknowledgment.
+Portions of the text in this manuscript were drafted with the assistance of large language models (LLMs), including ChatGPT [27]. All content was reviewed and verified by the authors.
 5 Data and Code Availability
-Our data and code can be found in https://github.com/SoroushVahidi/Ranking_with_MWFAS.
-As said before, our data is the datasets used in [20]; however, we have converted their files to .txt files that include an edges list, and the input of our program is the edge list of a graph, including edges and their weights.
+All data and source code used in this study are publicly available at: https://github.com/SoroushVahidi/Ranking_with_MWFAS.
+The datasets used in our experiments originate from [20] and were converted into text files containing weighted edge lists.
+Our implementation takes a weighted edge list as input, where each entry specifies a directed edge and its associated weight, and produces a ranking consistent with the acyclic structure induced by edge removals.